### v1

# Gaussian Herding across Pens:::: An Optimal Transport Perspective on Global Gaussian Reduction for 3DGS

Abstract
3D Gaussian Splatting (3DGS) has emerged as a powerful technique for radiance field rendering, but it typically requires millions of redundant Gaussian primitives, overwhelming memory and rendering budgets. Existing compaction approaches address this by pruning Gaussians based on heuristic importance scores, without global fidelity guarantee. To bridge this gap, we propose a novel optimal transport perspective that casts 3DGS compaction as global Gaussian mixture reduction. Specifically, we first minimize the composite transport divergence over a KD-tree partition to produce a compact geometric representation, and then decouple appearance from geometry by fine-tuning color and opacity attributes with far fewer Gaussian primitives. Experiments on benchmark datasets show that our method (i) yields negligible loss in rendering quality (PSNR, SSIM, LPIPS) compared to vanilla 3DGS with only 10% Gaussians; and (ii) consistently outperforms state-of-the-art 3DGS compaction techniques. Notably, our method is applicable to any stage of vanilla or accelerated 3DGS pipelines, providing an efficient and agnostic pathway to lightweight neural rendering.
Figure 1: Visual comparison. When reducing the number of Gaussians by 90%, our method outperforms other compaction techniques, such as PUP-3DGS, and remains competitive with the original 3DGS.
1 Introduction
Real-time 3D scene reconstruction and rendering dynamically generates photorealistic 3D representations from sensor data (e.g., multi-view images, LiDAR) with minimal latency, enabling critical applications in augmented/virtual reality (AR/VR), autonomous navigation, and immersive media [1, 2]. The current state-of-the-art, 3D Gaussian Splatting (3DGS) [3], iteratively learns 3D anisotropic Gaussian primitives with color and opacity attributes to model these scenes. During rendering, these 3D Gaussians are projected to 2D screens and $\alpha$-blended to achieve real-time photorealistic synthesis.
However, 3DGS faces significant efficiency challenges: its iterative densification process often produces millions of redundant Gaussians for complex scenes [4, 5, 6]. This inefficiency leads to high memory/storage costs and increased per-frame rendering time, limiting deployment on resource-constrained platforms like mobile and AR/VR devices [7].
Figure 2: Comparison of heuristic pruning and our method. The original mixture (left) with $10^{3}$ components is reduced to 5% using either pruning (middle) or our method (right). Our method better preserves the overall structure.
A common strategy to improve efficiency is compaction [8, 9, 10, 7]—reducing the number of Gaussians while preserving rendering fidelity. With fewer primitives, both storage requirements and rendering efficiency (particularly per-frame rendering time) are naturally improved. This approach is viable because 3DGS densification inherently generates redundant primitives [4, 5, 6]. Existing methods achieve compaction via pruning or random subset selection [11, 12, 5, 13, 14, 15, 6, 4, 16]. These strategies, which are based on naïvely discarding information, are simple to implement, but may be ineffective. In particular, they tend to lose critical structural details (as shown in Figure 1) or distort the underlying geometry (as shown in Figure 2). Such losses can degrade rendering quality, particularly in regions with fine-scale features or complex material properties. These limitations motivate our key question:
How to design an efficient compaction method that preserves 3D spatial and structural geometry?
To address this, we frame compaction as an optimization problem, where the goal is to approximate the original 3DGS representation with fewer Gaussians. Our solution leverages a statistical perspective, treating the scene’s geometric structure as a probabilistic model and employing principled reduction techniques to preserve fidelity.
  • (i)
    Geometric compaction via GMR. We first observe that the geometry of a 3DGS representation—defined by the positions, covariances, and opacities of its Gaussian primitives—can be interpreted as a Gaussian mixture model (GMM). Here, the mixture density is a convex combination of individual Gaussian densities, weighted by their opacities. This formulation naturally connects 3DGS compaction to Gaussian Mixture Reduction (GMR), a well-studied problem in statistics where a high-order GMM is approximated by one with fewer components while minimizing a divergence measure.
  • (ii)
    Appearance optimization. While geometry is compacted via GMR, the appearance of the scene—governed by the color attributes of the Gaussians—must also be preserved. To achieve this, we decouple the optimization of geometry (position, covariance) and appearance (color, opacity). After GMR-based compaction, we fine-tune the color and opacity of the reduced set of Gaussians using the standard 3DGS training pipeline. This two-stage strategy ensures that the compacted model maintains both geometric accuracy and photorealistic rendering quality.
For GMR, we minimize the composite transportation divergence [17], which is rooted in optimal transport theory [18, 19] and allows an effective algorithm. We tailor this GMR algorithm (detailed in Section 3.2.1) so that it scales well in scenes like 3DGS with an extensive amount of Gaussians. Crucially, our GMR optimizer does not merely select a subset of existing Gaussians; instead, it creates new primitives that can dynamically adjust their positions and covariances to better approximate the underlying geometry (as shown in Figure 2). Our method is fundamentally algorithm-agnostic: it functions as a plug-and-play module that can enhance both the standard 3DGS pipeline and any of its variants (Section 4), applicable at any stage of training to boost computational efficiency.
To summarize, our contributions are:
  • We open a new pathway to view 3DGS representations as a Gaussian mixture and perform compaction from the perspective of Gaussian mixture reduction via optimal transport.
  • We decouple geometry and appearance optimization in a scalable blockwise optimization scheme.
  • Empirical results demonstrate that our method preserves rendering quality using only 10% Gaussians.
  • We provide a plug-and-play module compatible with any 3DGS variant at any training stage.
2 Related Works
Compaction Techniques in 3DGS 3DGS compaction seeks to minimize the Gaussian count while preserving image quality. Existing work falls into two operations: densification (where to add) and pruning (what to drop). Most current approaches rely on per-Gaussian heuristic scores.
For densification, Taming-3DGS [10] ranks candidate Gaussians by combining gradient, pixel coverage, per-view saliency, and core attributes including opacity, depth, and scale; Color-cued GS [20] considers the view-independent spherical harmonics coefficient gradient to better capture color cues; and GaussianPro [9] guides growth using depth and normal maps. In the pruning phase, LightGaussian [11], Mini-Splatting [12], RadSplat [5], and AtomGS [13] compute an importance score from each Gaussian’s accumulated ray contribution, typically a mix of volume, opacity, transmittance, and hit count, and discard the lowest-ranked Gaussians. Gradient-aware variants prune by per-Gaussian gradients (Trimming-the-Fat) [14] or the second-order sensitivity score derived from the Hessian matrix (PUP-3DGS) [15]. The score can also be trained via a learnable mask, as in LP-3DGS [6], Compact3DGS [4], and HAC [16]. Moreover, multi-view consistency criteria discard Gaussians unseen by keyframes [21] or visible only in real but not virtual views [22]. For a comprehensive survey of 3DGS compression and compaction, we refer readers to [8, 23].
Despite their success, most existing strategies evaluate each Gaussian independently, leaving open the question of whether the retained set is truly the best global surrogate. Our work addresses this gap from a probabilistic perspective via Gaussian mixture reduction.
3DGS from Probabilistic Distribution Point of View Kheradmand et al. [24] formulate 3DGS as a Markov chain Monte Carlo process and use stochastic gradient Langevin dynamics to migrate dropped Gaussians onto retained ones, partially recycling lost information. However, their update remains pairwise and lacks convergence guarantees under a principled divergence. Moreover, this scheme is coupled with the original 3DGS pipeline, limiting its generalizability to other variants.
3 Method: Gaussian Herding across Pens
Figure 3: An illustration of the proposed GHAP-3DGS approach. The process begins with full-resolution 3DGS training to obtain initial geometric and appearance features. These Gaussians are then spatially partitioned using a KD-tree and grouped into blocks—analogous to sheep pens. We then perform blockwise Gaussian Model Reduction (GMR) to generate a small number of representative Gaussians within each block. This step is analogous to the popular kernel herding method  [25]. Finally, a lightweight appearance refinement step further optimizes the appearance feature of the reduced set. This multi-stage pipeline progressively guides the Gaussians in each block—analogous to herding across pens—toward a compact and high-fidelity representation.
We present a principled approach for compacting 3DGS representations through Gaussian Mixture Reduction (GMR). Our key insight is that The 3DGS densification strategy, while effective for view synthesis, produces redundant Gaussian components that can be consolidated without sacrificing visual quality. We achieve this by interpreting the 3D scene surface as an unnormalized probability density function (PDF), where the trained 3DGS representation serves as a Gaussian mixture approximation of this PDF. Thus, our problem reduces to finding a Gaussian mixture with fewer components to approximate the original, redundant one. Our procedure separately optimizing geometric (position $\mu$, covariance $\Sigma$) and appearance (opacity $\alpha$, color $c$) attributes to ensure rendering fidelity. The key idea of our method is illustrated in Figure 3.
3.1 Probabilistic Scene Representation
Let $\phi(x;\mu,\Sigma)=|2\pi\Sigma|^{-1}\exp(-(x-\mu)^{\top}\Sigma^{-1}(x-\mu))$ be the PDF of a Gaussian distribution with mean $\mu$ and covariance $\Sigma$. A Gaussian mixture with $n$ components is a distribution with density:
$\phi_{n}(x)=\sum_{i=1}^{n}\alpha_{i}\phi(x;\mu_{i},\Sigma_{i}),$
[…]
$\phi_{n}(x)\approx f(x),\quad\forall x\in{\mathcal{X}},$
where ${\mathcal{X}}$ is the 3D scene volume. Therefore, the geometry of the 3D scene can be effectively represented by a Gaussian mixture. Then, each of these Gaussian primitive is associated with its own color $c_{i}$. Both the geometry and appearance attributes are important for high quality rendering.
3.2 Compaction via Optimal Transport
Motivated by the observation that many 3DGS algorithms [4, 5, 6] produce a significant number of redundant Gaussians during training, we improve rendering efficiency through compaction—reducing the number of Gaussian primitives to achieve lower memory usage and faster rendering while preserving visual fidelity. The process consists of two key phases:
  1. 1.
    Geometric Compaction via GMR: Leveraging our probabilistic interpretation, we formulate compaction as Gaussian Mixture Reduction (GMR) [26], approximating the original Gaussian mixture with redundant components with fewer components. This yields a compacted geometric representation: $\bar{\phi}_{m}(x)=\sum_{j=1}^{m}\bar{\alpha}_{j}\phi(x;\bar{\mu}_{j},\bar{\Sigma}_{j})$, where $m\ll n$. This step modifies only the Gaussian positions ($\bar{\mu}_{j}$) and covariances ($\bar{\Sigma}_{j}$), leaving appearance attributes unchanged.
  2. 2.
    Appearance Optimization: The reduced Gaussians are initialized with appearance attributes (colors, opacities) and fine-tuned for optimal rendering performance. This step optimizes appearance only, maintaining geometric consistency.
Our approach decouples geometry and appearance optimization while using standard 3DGS training to preserve quality. Our training pipeline is visualized in Figure 3 and we describe the details below.
3.2.1 Geometric Compaction via GMR
Following Zhang et al. [17], we formulate compaction as minimizing the composite transportation divergence (CTD) between two Gaussian mixtures:
Definition 1 (Composite transportation divergence).
Let $c(\cdot,\cdot)$ be a divergence between two Gaussians. The composite transportation divergence (CTD) between two Gaussian mixtures $\phi_{n}(x)=\sum_{i=1}^{n}\alpha_{i}\phi(x;\mu_{i},\Sigma_{i})$ and $\phi^{\prime}_{m}(x)=\sum_{j=1}^{m}\alpha^{\prime}_{j}\phi(x;\mu^{\prime}_{j},\Sigma^{\prime}_{j})$ with cost function $c(\cdot,\cdot)$ is
${{\mathcal{T}}}_{c}(\phi_{n},\phi^{\prime}_{m})=\inf\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}\pi_{ij}c(\phi(\cdot;\mu_{i},\Sigma_{i}),\phi(\cdot;\mu^{\prime}_{j},\Sigma^{\prime}_{j})):\sum_{j=1}^{m}\pi_{ij}=\alpha_{i},\sum_{i=1}^{n}\pi_{ij}=\alpha^{\prime}_{j}\right\}.$
The CTD generalizes optimal transport [18] to mixtures, treating each component as a discrete distribution in Gaussian space. The cost function measures the cost of moving one unit of Gaussian from one location to another, and $\pi_{ij}$ measures the corresponding amount of mass that is being moved. The total cost is proportional to the cost and the mass, and the divergence is the smallest transportation cost to move the original mixture to the target mixture. The reduced mixture becomes
$\{\bar{\alpha}_{j},\bar{\mu}_{j},\bar{\Sigma}_{j}\}=\operatorname*{arg\,min}_{\{\alpha^{\prime}_{j},\mu^{\prime}_{j},\Sigma^{\prime}_{j}\}}{{\mathcal{T}}}_{c}(\phi_{n},\phi^{\prime}_{m}).$ (1)
With this formulation, the Gaussian mixture after compaction has optimal guarantee. The solution also guides the choice of $m$ to balance compactness and fidelity.
Algorithm 1 GMR via $k$-means Clustering
1:Initialize $\{\bar{\mu}_{j}^{(0)},\bar{\Sigma}_{j}^{(0)}\}_{j=1}^{m}$
2:for t=1,…, do
3:     Assignment Step:
4:     for $i=1$ to $n$ do
5:         Assign component $i$ to cluster ${\mathcal{C}}_{j}$ that minimizes $c(\phi(\cdot;\mu_{i},\Sigma_{i}),\phi(\cdot;\bar{\mu}_{j}^{(t-1)},\bar{\Sigma}_{j}^{(t-1)}))$
6:     end for
7:     Update Step:
8:     for $j=1$ to $m$ do
9:         Compute new cluster center: $\bar{\mu}_{j}^{(t)},\bar{\Sigma}_{j}^{(t)}=\operatorname*{arg\,min}\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}c(\phi(\cdot,\mu_{i},\Sigma_{i}),\phi(\cdot,\mu,\Sigma))$
10:     end for
11:     if no change in assignments then
12:         for $j=1$ to $m$ do
13:              Compute weight: $\bar{\alpha}_{j}=\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}$
14:         end for
15:         break
16:     end if
17:end for
As shown in Zhang et al. [17], (1) can be solved using the effective iterative algorithm in Algorithm 1. The algorithm reduces to a $k$-means variant in Gaussian space: 1) The assignment step follows the same principle as traditional $k$-means, but replaces the $L^{2}$ distance between vectors with a cost function $c(\cdot,\cdot)$ between Gaussian distributions. 2) The update step generalizes the cluster center computation: In traditional $k$-means, centers are updated as arithmetic averages (barycenters w.r.t. $L^{2}$ distance) of vectors in each cluster. In this algorithm, centers become barycenters of Gaussians in each cluster, minimized w.r.t. cost function $c(\cdot,\cdot)$. Thus, standard $k$-means emerges as a special case when using $L^{2}$ distance on vectorized Gaussian parameters.
While the standard GMR algorithm provides optimal theoretical guarantees, its direct application to 3DGS compaction proves computationally prohibitive. Although the algorithm must converge in finite steps [17] [^2] In particular, we find that in our experiment, the algorithm converges in only around 6 iterations., the assignment step involves $nm$ evaluations of the cost function per iteration. In typical 3D scenes, the number of Gaussians scales as $n=\Omega(10^{5})$ [^3] The $a_{n}=\Omega(b_{n})$ ($O(b_{n})$) if there exist $C\geq 0$ such that $a_{n}\geq Cb_{n}$ ($a_{n}\leq Cb_{n}$)., and even after 95% reduction, each iteration would still require at least $10^{8}$ operations and memory storage. For the update step, the computational cost for the cluster center depends on the pre-specified cost function $c(\cdot,\cdot)$. The KL divergence considered in Zhang et al. [17] suffers from its key limitations: it requires computing inverse covariance matrices, which introduces (1) numerical instability due to small eigenvalues of covariance matrices, and (2) significant overhead despite operating in ${\mathbb{R}}^{3}$ (with $O(\rho s^{2}\log n)$ inversions needed). To overcome this challenge, we introduce two key optimizations designed for 3DGS:
  • Blockwise GMR via KD-Tree: To improve computational and memory efficiency during training, we partition the scene into spatially localized blocks and perform GMR independently within each block. As demonstrated in Remark 1, this blockwise approach yields significant computational savings. While both KD-trees [27] and Octrees [28] are effective for spatial partitioning in 3D space, we employ a KD-tree for two key advantages. First, it produces more balanced partitions across regions. Second, it avoids unnecessary subdivisions in sparse regions that would waste computational resources.
    Our KD-tree is constructed solely from Gaussian centers $\{\mu_{i}\}_{i=1}^{n}$ (justified by the observed small eigenvalues of covariance matrices). Each split uses the median coordinate value, creating $2^{d}$ blocks at depth $d$. We set $d=\lfloor\log_{2}(n/s)\rfloor$ to ensure blocks contain at most $s$ Gaussians with $s\ll n$, then reduce each block to $m=\rho s$ components ($\rho$ = retention ratio).
    Remark 1 (Computational cost comparison).
    Our blockwise approach reduces the per-iteration computational cost from $O(\rho n^{2})$ to $O(\rho s^{2})$ per block. [^4] Each block reduces from $s$ Gaussian components to $\rho s$. With $2^{\text{depth}}=O(\log n)$ blocks in total, the overall complexity becomes $O(\rho s^{2}\log n)$. For typical values of $n=10^{5}$, $s=10^{3}$ and $\rho=0.05$, this reduces the cost from $10^{8}$ to approximately $10^{5}$ operations–a substantial improvement. The savings become even more pronounced for larger $n$. Furthermore, the reduction steps can be executed in parallel across blocks, offering additional computational speedup.
  • Efficient Cost Function: We propose a novel cost function that is both computationally efficient and maintains approximation quality, overcoming limitations of KL divergence used in [17]. We propose to use the following divergence:
    $c(\phi(\cdot;\mu,\Sigma),\phi(\cdot;\mu^{\prime},\Sigma^{\prime}))=\|\mu-\mu^{\prime}\|_{2}^{2}+\|\Sigma-\Sigma^{\prime}\|_{F}^{2},$ (2)
    which has three key advantages: First, it preserves distributional similarity as Gaussians are uniquely determined by their mean and covariance. Second, the assignment step requires only efficient vector/matrix norm computations. Third, the update step simplifies to weighted averages:
    $\bar{\mu}_{j}^{(t)}=\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}\mu_{i}/\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i},\quad\quad\quad\bar{\Sigma}_{j}^{(t)}=\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}\Sigma_{i}/\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}.$
3.2.2 Appearance Optimization
Algorithm 2 GHAP: 3DGS Compaction via Blockwise GMR
1:Input: Trained 3DGS model for $T$ steps to obtain $\{(\alpha_{i},\mu_{i},\Sigma_{i},c_{i})\}_{i=1}^{n}$, retention ratio $\rho$
2:Output: Compacted $\{(\bar{\alpha}_{j},\bar{\mu}_{j},\bar{\Sigma}_{j},\bar{c}_{j})\}_{j=1}^{m}$
3:Stage 1: Geometric Compaction
4:1. Build KD-tree from Gaussian centers $\{\mu_{i}\}_{i=1}^{n}$ with depth $d=\lfloor\log_{2}(n/s)\rfloor$
5:2. For each leaf block ${\mathcal{B}}_{k}$:
6: Run Algorithm 1 to reduce to $\rho s$ Gaussians
7:Stage 2: Appearance Optimization
8:1. Initialize appearance for each $\bar{\phi}_{j}$:
9:$\bar{c}_{j}\leftarrow c_{i^{*}}$ and $\bar{\alpha}_{j}\leftarrow\alpha_{i^{*}}$ where $i^{*}=\operatorname*{arg\,min}_{i\in[n]}\|\mu_{i}-\bar{\mu}_{j}\|_{2}$
10:2. Fine-tune $\{\bar{\alpha}_{j},\bar{c}_{j}\}$ using standard 3DGS rendering pipeline for $T$ steps
Following geometric compaction, we initialize the appearance attributes (opacity and colors) of the compacted Gaussian primitives. For each primitive in the reduced mixture, we assign the appearance parameters from its closest counterpart in the original Gaussian mixture. Using these initial values, we then optimize the appearance attributes through backpropagation following the standard 3DGS training pipeline. We optimize the opacity instead of directly using the opacity from the GMR algorithm since these output weights not necessarily satisfy the constraint on the opacity that it must be within 00 and $1$. Finetuning the the opacity can achieve better visualization performance.
3.3 Training Details with GHAP Algorithm
Integrating these components, we present our complete training pipeline in Algorithm 2, called Gaussian Herding Across Pens (GHAP). The process begins with standard 3DGS optimization for $T$ steps, followed by blockwise GMR. We then freeze the geometric parameters ($\mu$ and $\Sigma$) while fine-tuning appearance attributes (opacity $\alpha$ and color $c$) through an additional $T$-step 3DGS optimization. This procedure can be iteratively applied throughout training as needed.
4 Experiments
4.1 Experimental Setup
Datasets
For a comprehensive evaluation of our GHAP algorithm, we use three real-world datasets: Tanks and Temples [29], Mip-NeRF 360 [30], and Deep Blending [31], which cover varying levels of detail, lighting conditions, and scene complexities. For each dataset, we adopt the same scenes as in [8].
  • Tanks and Temples: We evaluate two unbounded outdoor scenes, “Truck” and “Train”, both featuring centered viewpoints.
  • Mip-NeRF 360: We test on a mix of indoor and outdoor scenes, including “Bicycle”, “Bonsai”, “Counter”, “Flowers”, “Garden”, “Kitchen”, “Room”, “Stump”, and “Treehill”, all with centered viewpoints.
  • Deep Blending: We include two indoor scenes, “Dr. Johnson” and “Playroom”, where the viewpoint is directed outward.
Baselines
To evaluate the effectiveness of our proposed method, we compare it against several strong baselines in the domain of 3DGS compaction: Mini-Splatting-D [12], AtomGS [13], LightGaussian [11], and PUP-3DGS [15]. We set $s=3\times 10^{3}$. The sensitivity analysis of $s$ is given in Appendix. The initialization of the iterative algorithm in (1) randomly selects $m$ Gaussians per block.
Evaluation Metrics We assess 3DGS compaction using standard metrics for rendering quality and efficiency. For quality, we report: (1) PSNR, measuring pixel-level accuracy; (2) SSIM, evaluating perceptual similarity based on luminance, contrast, and structure; and (3) LPIPS, capturing perceptual distance via a learned model. Higher PSNR/SSIM and lower LPIPS indicate better quality. For efficiency, we measure FPS across retention ratios, where lower retention yielding higher FPS implies greater rendering efficiency.
Device
All experiments were conducted on a server with 256 GB RAM and a 96-core Intel® Xeon® Platinum 8255C CPU, and on a workstation equipped with five NVIDIA RTX 3090 GPUs, each with 24 GB of VRAM.
4.2 Quantitative Results
We first demonstrate the effectiveness of our GHAP compaction method as a plug-in module within 3DGS and its variants. Second, we show that our approach outperforms other SOTA compaction techniques. We present the results first and summarize our findings at the end of the section.
As a Plug-In Compaction Method
Our method can be used as a plug-in compaction method in various 3DGS training algorithms. This demonstrates the broad applicability of our proposed method. To verify this, we apply our compaction method within various 3DGS pipelines. Specifically, we consider three representative variants: 3DGS-30K, AtomGS, and Mini-Splatting-D. Each of them employs a distinct densification strategy, and our method can be directly embedded into the pipeline without extensive engineering effort.
For each variant, we evaluate performance under two retention ratios (10% and 20%) and report results using three standard metrics: PSNR, SSIM, and LPIPS. Table 1 summarizes our experimental results.
Table 1: Quantitative evaluation of our GHAP method with respect to compaction performance applied to different 3DGS variants under varying retention ratios (10% and 20%).
Method $\rho$ Tanks&Temples MipNeRF-360 Deep Blendering
[…]
3DGS-30k 10% 0.818 23.312 0.242 0.764 26.404 0.314 0.905 29.647 0.264
20% 0.835 23.615 0.212 0.788 26.973 0.275 0.907 29.864 0.252
100% 0.853 23.785 0.169 0.813 27.554 0.221 0.907 29.816 0.238
Mini-Splatting-D 10% 0.835 23.232 0.198 0.802 27.090 0.250 0.909 30.042 0.254
20% 0.855 23.403 0.171 0.821 27.310 0.214 0.912 30.170 0.238
100% 0.848 23.338 0.140 0.832 27.486 0.176 0.907 29.980 0.211
AtomGS 10% 0.793 22.988 0.274 0.764 26.535 0.307 0.899 29.347 0.282
20% 0.812 23.282 0.240 0.788 27.025 0.269 0.902 29.347 0.268
100% 0.814 23.289 0.235 0.796 27.135 0.251 0.902 29.314 0.267
Comparison with SOTA Compaction Methods
We compare our method with two recent compaction baselines: PUP-3DGS and LightGaussian. To ensure a fair evaluation, all methods are built upon the same backbone model, 3DGS-30K. Compaction is performed at the same stage, when the number of Gaussians reaches 15K, followed by 15K fine-tuning iterations. During fine-tuning, we use identical optimization settings across all methods, including learning rate and other hyperparameters. Two compaction ratios are considered: 10% and 20%. A detailed comparison of rendering quality across different compaction methods and retention ratios is presented in Table 2.
Table 2: We compare our GHAP method against LightGaussian and PUP-3DGS, modifying their original workflows to align with our compaction pipeline.
Method-Compact Rate Tandt&Temples MipNeRF-360 Deep Blendering
[…]
3DGS-30k-100% 0.853 23.785 0.169 0.813 27.554 0.221 0.907 29.816 0.238
GHAP(Ours)-10% 0.818 23.312 0.242 0.764 26.404 0.314 0.905 29.647 0.264
Light-GS-10% 0.791 22.560 0.248 0.779 26.585 0.261 0.861 27.136 0.326
Pup-GS-10% 0.767 21.519 0.28 0.753 25.332 0.309 0.895 29.153 0.274
GHAP(Ours)-20% 0.835 23.615 0.212 0.788 26.973 0.275 0.907 29.864 0.252
Light-GS-20% 0.791 22.559 0.248 0.779 26.588 0.261 0.861 27.142 0.326
Pup-GS-20% 0.809 22.603 0.228 0.790 26.671 0.257 0.905 29.719 0.248
Figure 4: Visualization before and after compaction. Various scenes before and after compaction under different 3DGS algorithms. Our GHAP method yields negligible loss in rendering quality.
Takeaways
Based on these experiments, we draw four key observations:
  • Superior fidelity retention. Our method substantially outperforms both LightGaussian and PUP-3DGS in SSIM and PSNR. Surprisingly, our method can even achieve better performance than the original method, this is demonstrated in Table 1 on the Deep Blending scene with the Mini-Splatting-D backbone and at a 10% retention ratio. Likewise, in Table 2, we consistently surpass LightGaussian and PUP-3DGS in most datasets and metrics.
  • Efficiency gains. At a 10% Gaussian primitives retention rate, our compaction framework yields a 140.3% increase in rendering speed (FPS), demonstrating that significant model slimming can be achieved without sacrificing and even often improving performance (see ablation study).
  • Cross-dataset consistency. Performance remains stable across datasets. In outdoor scenes (Tanks and Temples), SSIM value fidelity drops ratio (relative drop in rendering metrics after compaction) less than 4.1%. Meanwhile, for the indoor scene, the PSNR value retains 99.4% of the original value.
  • LPIPS sensitivity. We observe that our GMR compaction method incurs a relatively larger degradation in LPIPS compared to SSIM and PSNR, although the reduction remains within acceptable bounds. This trend aligns with prior findings [11, 15], as LPIPS is more sensitive to structural and textural distortions. Since compaction inherently reduces model capacity, fine-grained texture reconstruction becomes more challenging, which explains the behavior observed in our qualitative and quantitative analyses
Qualitative Visualization
We visualize two scenes for each baseline 3DGS algorithm before and after compaction in Figure 4. Across most scenes, our compaction method successfully preserves rendering quality using only 10% of the Gaussian primitives. By design, our reduced Gaussians approximate the original Gaussians from the 3DGS algorithm. If the original model lacks good visual quality due to lack to Gaussian primitives in certain regions, our method cannot perform well—as demonstrated in the ‘Bicycle’ scene, where the poor performance of the Atom-GS baseline prevents our approach from maintaining high fidelity after compaction. Interestingly, in some cases, our post-compaction results surpass the pre-compaction quality. A representative example is the ‘Kitchen’ scene based on Mini-Splatting-D. We conjecture this occurs because Mini-Splatting-D lacks a pruning mechanism, often generating an oversaturated set of Gaussians that introduce visual artifacts (e.g., unnatural shadows in the lower-left region). Our method mitigates this issue by globally regularizing the Gaussian distribution: improving expressiveness, reducing unnecessary density, and preserving the underlying 3D structure.
4.3 Ablation Studies
GMR versus Random Subsampling We evaluate the contribution of each stage in our pipeline—geometric compaction and appearance optimization—on the Tanks and Temples dataset. For comparison, we use random subsampling at the same compaction ratio. This allows us to assess the effectiveness of our design choices at each stage. The experiments show that our compaction method is significantly more effective than random subsampling,
Table 3: Ablation study. Mean PSNR, SSIM, and LPIPS on the Tanks and Temples dataset at each stage of GHAP pipeline, with random subsampling as a control. Each operation is applied cumulatively to all subsequent stages.
Method Tanks&Temples
[…]
3DGS-15K 0.839 23.084 0.194
+ 10% GHAP Compaction 0.502 14.015 0.483
+ 10% Random Compaction 0.333 9.573 0.555
+15K Fine-tuning 0.818 23.312 0.242
+15K Fine-tuning 0.712 21.312 0.282
producing a more meaningful scene representation. Appearance optimization further leads to substantial quantitative improvements. Together, these results validate the necessity and effectiveness of both stages in our proposed pipeline.
Figure 5: GHAP efficiency-quality trade-off. SSIM and FPS (top) and LPIPS and PSNR (bottom) across different retention ratios $\rho$.
Rendering Metrics under GHAP
To analyze the efficiency-quality trade-off of the proposed GHAP approach, we vary the retention ratio $\rho$ from $10\%$ to $100\%$. A lower retention ratio enables more aggressive compaction, resulting in faster rendering speed but potentially compromising rendering quality. Our results demonstrate that decreasing the retention ratio yields 2×2$\times$2 × faster rendering with only marginal quality degradation. For instance, as $\rho$ decreases from $100\%$ to $10\%$: LPIPS increases modestly from 0.14 to 0.21; SSIM declines slightly from 0.885 to 0.860; PSNR remains stable between 25.0 dB and 25.5 dB. These findings indicate that retaining even a modest number of Gaussians effectively preserves visual quality, achieving an excellent performance-fidelity trade-off.
5 Conclusion
We present a novel optimal transport-based global Gaussian reduction approach for 3D Gaussian Splatting, addressing the inefficiencies of existing Gaussian pruning and compaction methods. By formulating the problem as Gaussian mixture reduction, our method minimizes composite transport divergence to derive a compact geometric representation, followed by appearance fine-tuning with significantly fewer primitives. It provides a theoretically sound foundation for efficient neural rendering. Extensive experiments demonstrate that our approach achieves rendering quality nearly indistinguishable from vanilla 3DGS while retaining only  10% of the Gaussians, outperforming existing compaction techniques. Numerical results show our framework is versatile, compatible with various 3DGS pipelines. This work advances the practical deployment of 3DGS by striking an optimal balance between computational efficiency and visual fidelity. Future work includes extending our method to dynamic 3DGS for real-time temporal scenes and developing adaptive reduction strategies for task-specific needs.
Acknowledgments and Disclosure of Funding
Qiong Zhang is supported by the National Natural Science Foundation of China Grant 12301391.
Appendix
Appendix A More Details about the Blockwise GMR Compaction
A.1 Introduction to Optimal Transport (OT)
The Optimal Transport (OT) problem dates back to 1781, when French mathematician Gaspard Monge (1746–1818) first formulated it as finding the minimal-cost way to move sand into a hole. This “transport plan” minimizes the transportation cost, hence the term “optimal transport”. The original formulation is known as Monge’s problem. In 1947, Russian economist Leonid Kantorovich (1912–1986) relaxed Monge’s formulation, leading to the so-called Monge-Kantorovich problem.
Specifically, let $P$ and $Q$ be two distributions over a metric space ${\mathcal{X}}$, let $c:{\mathcal{X}}\times{\mathcal{X}}\to{\mathbb{R}}_{+}$ be a cost function, and let the coupling
$\Pi(P,Q)=\left\{\pi(x,y):\int\pi(dx,\cdot)=Q(\cdot),\int\pi(\cdot,dy)=P(\cdot)\right\}$
be the set of joint distributions with marginals $P$ and $Q$. For any cost function $c$, the total transportation cost induced by a plan $\pi\in\Pi(P,Q)$ is defined as
${\mathcal{I}}_{c}(\pi)=\int_{\mathcal{X}\times\mathcal{X}}c(x,y)\,\pi(dx,dy).$
Here, $\pi(x,y)$ indicates how much “mass” is transported from $x$ to $y$. The first constraint, $\int\pi(x,dy)=P(x)$, ensures that the mass at location $x$ is spread over ${\mathcal{X}}$, while the second constraint, $\int\pi(dx,y)=Q(y)$, ensures that the destination at $y$ receives the required mass. The OT problem seeks the optimal plan
$\pi^{*}=\operatorname*{arg\,min}\left\{{\mathcal{I}}_{c}(\pi):\pi\in\Pi(P,Q)\right\},$
minimizing the transportation cost. The corresponding minimal cost,
$\mathcal{T}_{c}(P,Q)={\mathcal{I}}_{c}(\pi^{*}),$
induces a divergence between $P$ and $Q$. This OT divergence provides a principled way to measure distributional similarity, enabling applications in density matching, generative modeling, and dimensionality reduction.
The optimal plan $\pi^{*}$ and OT divergence typically lack closed-form solutions. Numerical algorithms [18] are used to compute the OT between discrete measures. When $P$ and $Q$ are discrete, the OT divergence reduces to:
Example 1 (OT Divergence Between Discrete Distributions).
For $P=\sum_{i=1}^{n}u_{n}\delta_{x_{n}}$ and $Q=\sum_{j=1}^{m}v_{m}\delta_{y_{m}}$, the OT divergence is
${\mathcal{T}}_{c}(P,Q)=\min\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}\pi_{ij}c(x_{i},y_{j}):\sum_{i=1}^{n}\pi_{ij}=v_{j},\sum_{j=1}^{m}\pi_{ij}=u_{i}\right\}.$ (3)
Exact solutions can be found via linear programming, while approximations are obtained using algorithms like Sinkhorn.
A.2 Relationship Between Composite Transportation Divergence and OT
The Composite Transportation Divergence (CTD) [32, 33] extends OT to Gaussian mixtures. For two mixtures $\phi_{n}=\sum_{i=1}^{n}\alpha_{i}\phi(\cdot;\mu_{i},\Sigma_{i})$ and $\phi^{\prime}_{m}=\sum_{j=1}^{m}\alpha^{\prime}_{j}\phi(\cdot;\mu^{\prime}_{j},\Sigma^{\prime}_{j})$, the CTD is:
${{\mathcal{T}}}_{c}(\phi_{n},\phi^{\prime}_{m})=\inf\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}\pi_{ij}c(\phi(\cdot;\mu_{i},\Sigma_{i}),\phi(\cdot;\mu^{\prime}_{j},\Sigma^{\prime}_{j})):\sum_{j=1}^{m}\pi_{ij}=\alpha_{i},\sum_{i=1}^{n}\pi_{ij}=\alpha^{\prime}_{j}\right\},$
Comparing this with the discrete OT divergence (3), the CTD treats Gaussian mixtures as discrete distributions over the space of Gaussians, defining the divergence as the OT between them.
To illustrate, consider $n$ warehouses and $m$ factories in the space of Gaussian distributions ${\mathcal{F}}$. The $i$th warehouse, at $\phi(\cdot;\mu_{i},\Sigma_{i})$, holds $\alpha_{i}$ units of material, while the $j$th factory, at $\phi(\cdot;\mu^{\prime}_{j},\Sigma^{\prime}_{j})$, requires $\alpha^{\prime}_{j}$ units. The cost to transport material from $i$ to $j$ is $c(\phi(\cdot;\mu_{i},\Sigma_{i}),\phi(\cdot;\mu^{\prime}_{j},\Sigma^{\prime}_{j}))$, and $\pi_{ij}\geq 0$ denotes the transported amount. The total cost under plan $\pi$ is $\sum_{i,j}\pi_{ij}c(\phi(\cdot;\mu_{i},\Sigma_{i}),\phi(\cdot;\mu^{\prime}_{j},\Sigma^{\prime}_{j}))$. The coupling set $\Pi(\alpha,\alpha^{\prime})=\{\pi_{ij}:\sum_{j=1}^{m}\pi_{ij}=\alpha_{i},\sum_{i=1}^{n}\pi_{ij}=\alpha^{\prime}_{j}\}$ ensures: (a) correct material removal from warehouses, and (b) correct delivery to factories. The OT problem seeks the plan $\pi^{*}\in\Pi(\alpha,\alpha^{\prime})$ minimizing the total cost. The minimal cost corresponds to the CTD between the mixtures, quantifying the optimal transport cost between them.
Our compaction method leverages this interpretation, approximating a mixture with fewer Gaussians by minimizing their CTD-based dissimilarity.
A.3 Algorithm Intuition
At first glance, the optimization problem in (1) is bilevel: the OT plan must be found for any candidate $\{\bar{\alpha}_{j},\bar{\mu}_{j},\bar{\Sigma}_{j}\}_{j=1}^{m}$, and the objective must be minimized. However, as shown by Zhang et al. [17], this simplifies with a clear interpretation. The column-wise marginal constraints on $\pi$ are redundant, and each Gaussian primitive $(\mu_{i},\Sigma_{i})$ transports its full mass to its “closest” counterpart. The optimal plan thus corresponds to clustering $\{(\mu_{i},\Sigma_{i})\}_{i=1}^{n}$ into $m$ clusters $\{{\mathcal{C}}_{j}\}_{j=1}^{m}$, with the original Gaussians forming cluster “barycenters”. Mathematically, the simplified program is:
$\min\left\{\sum_{j=1}^{m}\sum_{i\in{\mathcal{C}}_{j}}\pi_{ij}c(\phi(\cdot;\mu_{i},\Sigma_{i}),\phi(\cdot;\bar{\mu}_{j},\bar{\Sigma}_{j}))\colon[n]={\mathcal{C}}_{1}\sqcup\cdots\sqcup{\mathcal{C}}_{m}\right\},$
where $\sqcup$ denotes disjoint union. The optimal plan has a closed form:
$\pi_{ij}=\begin{cases}\alpha_{i}&\text{if }j=\operatorname*{arg\,min}_{k}c\left(\phi(\cdot;\mu_{i},\Sigma_{i}),\phi(\cdot;\bar{\mu}_{k},\bar{\Sigma}_{k})\right),\\ 0&\text{otherwise}.\end{cases}$
With this, $\pi$ and the reduced mixture parameters can be updated alternately, formalized in the following $k$-means-like algorithm:
  • Assignment Step: Each Gaussian $\phi(\cdot;\mu_{i},\Sigma_{i})$ is assigned to cluster $\mathcal{C}_{j}$ by minimizing $c(\phi(\cdot;\mu_{i},\Sigma_{i}),\phi(\cdot;\bar{\mu}_{j},\bar{\Sigma}_{j}))$, analogous to $k$-means’ nearest-centroid assignment.
  • Update Step: For the cost function
    $c(\phi,\phi^{\prime})=\|\mu-\bar{\mu}\|_{2}^{2}+\|\Sigma-\bar{\Sigma}\|_{F}^{2},$
    the parameters $(\bar{\mu}_{j},\bar{\Sigma}_{j})$ are updated as weighted averages:
    $\bar{\mu}_{j}=\frac{\sum_{i\in\mathcal{C}_{j}}\alpha_{i}\mu_{i}}{\sum_{i\in\mathcal{C}_{j}}\alpha_{i}},\quad\bar{\Sigma}_{j}=\frac{\sum_{i\in\mathcal{C}_{j}}\alpha_{i}\Sigma_{i}}{\sum_{i\in\mathcal{C}_{j}}\alpha_{i}}.$
    This mirrors $k$-means’ centroid update.
A.4 Algorithm Convergence
The following theorem guarantees convergence of the CTD sequence generated by Algorithm 1. In the worst case, the algorithm requires exponentially many steps, but it typically converges in about 5 iterations. A full proof is given in Zhang et al. [17].
Theorem 1 (Convergence of the Algorithm).
Suppose $c(\cdot,\cdot)$ is continuous, and for any $\Delta>0$ and $\phi^{\star}$, the set $\{\phi:c(\phi^{*},\phi)\leq\Delta\}$ is compact under some Gaussian space metric. Let $\{\bar{\phi}_{m}^{(t)}\}$ be the sequence generated by the update step with initial $\{\bar{\alpha}_{j}^{(0)},\bar{\mu}_{j}^{(0)},\bar{\Sigma}_{j}^{(0)}\}$, and let $\mathcal{T}_{c}^{(t+1)}$ be the CTD at iteration $t$. Then:
  1. 1.
    There exists $T$ and $\{\bar{\alpha}_{j}^{*},\bar{\mu}_{j}^{*},\bar{\Sigma}_{j}^{*}\}$ such that for all $t\geq T$, $\{\bar{\alpha}_{j}^{(t)},\bar{\mu}_{j}^{(t)},\bar{\Sigma}_{j}^{(t)}\}=\{\bar{\alpha}_{j}^{*},\bar{\mu}_{j}^{*},\bar{\Sigma}_{j}^{*}\}$ and $\mathcal{T}_{c}^{(t+1)}=\mathcal{T}_{c}^{(*)}$, where $\mathcal{T}_{c}^{(*)}$ is the CTD between the original mixture and $\{\bar{\alpha}_{j}^{*},\bar{\mu}_{j}^{*},\bar{\Sigma}_{j}^{*}\}$.
  2. 2.
    The limit point $\{\bar{\alpha}_{j}^{*},\bar{\mu}_{j}^{*},\bar{\Sigma}_{j}^{*}\}$ is a local minimum of ${\mathcal{T}}_{c}$.
  3. 3.
    An MM-based exhaustive algorithm with $O(m^{n})$ complexity solves (1).
Appendix B More Background and Involved Techniques in Paper
B.1 Gaussian splatting and efficient variants
Due to its advantages in rendering speed and visual fidelity, 3DGS has been adopted in a broad range of applications, including human reconstruction, AI-generated content, autonomous driving, and beyond [34, 35, 36, 37]. Extensions to dynamic 3DGS, editable 3DGS, and surface representation have further broadened its utility [38, 39, 40]. However, the standard 3DGS suffers from significant storage, memory, and rendering costs when applied to large-scale or complex scenes. This limits its deployment on resource-constrained platforms and in broader areas.
To improve storage and rendering efficiency while maintaining visual quality, recent research has focused on two complementary directions: compression and compaction [8]. Compression methods shrink file size by exploiting redundancy in Gaussian attributes (e.g., using vector quantization to cluster similar parameters) [11, 41] or structuring Gaussians into more compact forms (e.g., grids, anchors, and shared codes) [42, 43]. On the other hand, compaction aims to reduce the number of Gaussians required for accurate rendering by optimizing their spatial distribution and pruning redundant primitives [12, 6, 5]. Hybrid pipelines such as LightGaussian [11] and Octree-GS [7] combine compression and compaction together to deliver highly compact and lightweight neural renders.
B.2 Detailed 3DGS pipline
The advantages of 3DGS in rendering speed and image fidelity have made it applicable to a wide range of tasks, including human reconstruction, AI-generated content, autonomous driving, and beyond [34, 35, 36, 37]. Extensions to dynamic 3DGS, editable 3DGS, and surface representation have further broadened its utility [38, 39, 40]. 3D Gaussian Splatting (3DGS) represents a scene as a set of anisotropic 3D Gaussian primitives, each parameterized by its spatial location, shape, opacity, and radiance. The pipeline consists of the following key steps:
  1. 1.
    Initialization. From Structure-from-Motion (SfM), obtain calibrated camera poses and a sparse point cloud. Each point is initialized as a 3D Gaussian with an opacity $\alpha_{i}$:
    $\phi(x;\mu_{i},\Sigma_{i})=\left|2\pi\Sigma\right|^{-1}\exp\left(-\frac{1}{2}\left(x-\mu_{i}\right)^{T}\Sigma_{i}^{-1}\left(x-\mu_{i}\right)\right),$
    where $\mu_{i}\in\mathbb{R}^{3}$ is the position (mean), $\Sigma_{i}\in\mathbb{R}^{3\times 3}$ is the covariance matrix (anisotropic shape), and $\alpha_{i}\in[0,1)$.
  2. 2.
    Projection & Rasterization. Each 3D Gaussian is projected to 2D using the camera model,
    $\Sigma_{i}^{\prime}=J{W}\Sigma_{i}J{W}^{T},$
    where $W$ is a veiw transformation matrix and $J$ is the Jacobian of the projective transform. Rasterization is done using a differentiable splatting approach which makes optimization possible.
  3. 3.
    Image Formation (Alpha Blending). The pixel color is computed via volumetric blending:
    $C=\sum_{i=1}^{N}T_{i}\cdot\alpha_{i}\cdot c_{i}\text{ with }T_{i}=\prod_{j=1}^{i-1}\left(1-\alpha_{j}\right),$
    where $c_{i}$ is the SH-predicted color of the $i$-th Gaussian in front-to-back order.
  4. 4.
    Optimization. The Gaussian parameters $\left\{\mu_{i},\Sigma_{i},\alpha_{i},c_{i}\right\}$ are optimized to minimize a photometric loss:
    $\mathcal{L}=\left(1-\lambda\right)\left\|\widehat{\bm{C}}-\bm{C}^{\star}\right\|_{1}+\lambda\cdot\mathcal{L}_{\text{SSIM}},$
    where $\widehat{\bm{C}}$ is the rendered image, $\bm{C}^{\star}$ the ground truth, and $\lambda\in[0,1]$ balances the two loss terms.
  5. 5.
    Adaptive Densification & Pruning. During training, Gaussians are cloned (under-reconstruction) or split (over-reconstruction) based on view-space gradient magnitude, and low-contribution Gaussians are pruned.
B.3 Involved 3DGS Variants
Although 3DGS features a well-designed pipeline and an efficient differentiable rasterization scheme, its densification process still suffers from several limitations at the detail level. For example, the non-uniform distribution of Gaussians can lead to overlapping primitives in dense regions and under-reconstruction in sparse areas. In addition, the optimization tends to favor large Gaussians that quickly cover the scene to achieve fast convergence. This behavior, however, often results in poor reconstruction of high-frequency details and introduces geometric noise.
Consequently, a growing body of work has focused on refining the densification stage of the original 3DGS framework. In this section, we highlight two such methods—Atom-GS [13] and Mini-Splatting (-D) [12]—which specifically enhance the densification process. While these methods also modify other components such as pruning, we isolate and utilize only the early-stage point cloud generated before 15K training iterations. Applying our compaction method to these two distinct variants further illustrates its generality.
It is important to note that our approach is broadly applicable to any point cloud produced by 3DGS-style pipelines, as long as each Gaussian retains its mean, covariance, and opacity, along with a reasonable optimization strategy for opacity and color.
Atom-GS
To address the tendency of 3DGS to rely on large Gaussians for rapid scene coverage, AtomGS replaces the free-form Gaussian optimization in 3DGS with a structured formulation based on atom Gaussians—uniform, isotropic primitives that better adhere to the underlying scene geometry. In addition, AtomGS introduces geometry-guided optimization to further enhance the spatial fidelity of the reconstructed scene. AtomGS builds upon two key techniques:
  1. 1.
    Atomized Proliferation.
    • Constrains Gaussians in high-detail regions into fixed-size, isotropic Atom Gaussians.
    • Uses a geometric schedule to refine scale and enhance surface coverage over training iterations.
    • Merges Gaussians in flat areas to reduce redundancy.
  2. 2.
    Geometry-Guided Optimization.
    • Introduces Edge-Aware Normal Loss by weighting normal smoothness using an edge map derived from the RGB image.
    • Preserves sharp features while smoothing flat regions, improving alignment with the true 3D geometry.
Mini-Splatting (-D)
The original 3DGS often suffers from non-uniform Gaussian distributions, resulting in excessive overlap in dense regions and insufficient reconstruction in sparse areas. Mini-Splatting addresses this issue by reorganizing the spatial distribution of Gaussians through a two-stage pipeline consisting of densification and simplification. Its goal is to reduce the total number of Gaussians without compromising rendering quality. Mini-Splatting-D is a variant of Mini-Splatting, tailored for quality-oriented rendering. While it follows a similar overall pipeline, the key difference lies in the omission of the simplification stage, allowing the method to focus entirely on maximizing rendering quality. In our implementation, we focus solely on the densification process used in our pipeline.
  • Blur Split: Identifies Gaussians that dominate blurry regions (via screen-space analysis) and splits them to recover fine detail.
  • Depth Reinitialization: Reprojects depth maps derived from midpoints between ray-ellipsoid intersections to generate dense and uniform 3D points for reinitialization.
Beyond evaluating compatibility with different 3DGS densification variants, we compare our method against recent state-of-the-art compaction techniques. Specifically, we consider two representative baselines: LightGaussian [11] and PUP-3DGS [15]. A brief overview of each method is provided below.
LightGaussian
Original 3DGS overpopulates Gaussians during densification, causing storage bloat and rendering inefficiency. LightGaussian prunes redundant Gaussians based on a global significance score (computed via hit count, opacity, and normalized volume), followed by fine-tuning to recover lost details. The steps include:
  1. 1.
    Compute significance for each Gaussian by a score based on the volume of the Gaussian sphere and the number of covered pixels.
  2. 2.
    Prune low-score Gaussians.
  3. 3.
    Fine-tune retained Gaussians with photometric loss.
PUP-3DGS
LightGaussian employs a visibility-aware importance score, which tends to favor pruning small Gaussians while retaining larger background primitives. In contrast, PUP-3DGS introduces a mathematically principled importance measure based on second-order derivatives, allowing it to preserve small but geometrically critical Gaussians in foreground regions. The technical steps are:
  1. 1.
    Fisher Approximation. Computes the gradients of the Gaussian parameters to approximate the Hessian matrix, which is then used to derive the sensitivity score.
  2. 2.
    Iterative Pruning. Rank Gaussians by sensitivity scores, prune low-sensitivity ones, and fine-tune iteratively to recover lost fidelity.
Appendix C Sensitivity Analysis
C.1 Choice of KD-tree Depth
In our algorithm, the KD-tree depth $d$ (or equivalently, the number $s$ of Gaussians per KD block) is treated as a hyperparameter. We investigate how rendering quality varies with different values of $d$ on the Truck scene. The results are shown in Figure 6.
Figure 6: Sensitivity Analysis of $d$. Our GHAP algorithm exhibits stable rendering performance across varying KD-tree depths $d$.
We observe that, within a reasonable range, the KD-tree depth $d$ has minimal impact on rendering quality. In our experiments on the Truck scene, setting $d<6$ results in distance matrices that are too large to fit into GPU memory, causing memory overflow. Conversely, when $d>16$, the number of partitions becomes excessively large, and most blocks contain one or zero Gaussians, rendering the algorithm inapplicable.
C.2 Sampling Iteration
Since most 3DGS-based algorithms complete the densification stage before 15K training iterations, we apply our sampling procedure after this point to achieve a target retention ratio. To study the effect of sampling at different post-densification stages, we evaluate our method at various iterations beyond 15K. The experimental results are presented in Figure 7.
Figure 7: Sensitivity Analysis of Sampling Iteration. The rendering performance of our GHAP algorithm has negligible loss by variations in sampling iteration.
As shown in the results, sampling at any iteration beyond 15K has minimal impact on the final rendering quality. These results indicate that continuing optimization beyond 15K provides diminishing returns. Moreover, once sampling is performed, only a small number of fine-tuning iterations are needed for the compacted model to converge. Based on these observations, we recommend applying sampling uniformly at the 15K iteration, followed by an additional 15K fine-tuning iterations for all experiments.
Appendix D Additional Scene Visualizations
We report a more comprehensive set of results for the plug-in experiments in Table 4, including two additional metrics: the number of Gaussians (in thousands) and rendering speed measured in FPS. As shown in the updated results, our compaction method significantly improves FPS compared to the original models, often yielding over a 2$\times$ speedup in rendering performance.
In addition to the previously shown Figure 4, we present more detailed qualitative comparisons across multiple scenes in Figure 8. Our method consistently preserves the visual quality of the original models. Notably, when applied to stronger 3DGS variants with improved densification strategies—such as Mini-Splatting-D—our compaction framework performs even better. This is reflected in the fact that, after compaction, Mini-Splatting-D often achieves higher rendering quality than the original 3DGS baseline.
Table 4: Additional information in plug-in experiment.
Method-Compact Rate $\rho$ Tank&Temples MipNeRF-360 Deep Blendering
SSIM PSNR LPIPS $k$ Gaussians FPS SSIM PSNR LPIPS $k$ Gaussians FPS SSIM PSNR LPIPS $k$ Gaussians FPS
3DGS-30k 10% 0.818 23.312 0.242 157 341 0.764 26.404 0.314 272 272 0.905 29.647 0.264 248 290
20% 0.835 23.615 0.212 313 300 0.788 26.973 0.275 545 230 0.907 29.864 0.252 496 237
100% 0.853 23.785 0.169 1577 168 0.813 27.554 0.221 2625 126 0.907 29.816 0.238 2476 126
Mini-Splatting-D 10% 0.835 23.232 0.198 147 564 0.802 27.090 0.25 318 389 0.909 30.042 0.254 156 549
20% 0.855 23.403 0.171 293 526 0.821 27.310 0.214 636 299 0.912 30.17 0.238 318 489
100% 0.848 23.338 0.14 1473 209 0.832 27.486 0.176 3177 111 0.907 29.980 0.211 1597 197
Atom-GS 10% 0.793 22.988 0.274 426 401 0.764 26.535 0.307 468 304 0.899 29.347 0.282 464 419
20% 0.812 23.282 0.24 852 283 0.788 27.025 0.269 937 236 0.902 29.347 0.268 925 323
100% 0.814 23.289 0.235 4274 87 0.796 27.135 0.251 4699 86 0.902 29.314 0.267 4625 121
Figure 8: More scenes visualization. Visual comparisons across additional scenes before and after compaction.

### v2

@@
-Gaussian Herding across Pens:::: An Optimal Transport Perspective on Global Gaussian Reduction for 3DGS
+Gaussian Herding across Pens:: An Optimal Transport Perspective on Global Gaussian Reduction for 3DGS
 Abstract
@@
 Notably, our method is applicable to any stage of vanilla or accelerated 3DGS pipelines, providing an efficient and agnostic pathway to lightweight neural rendering.
+The code is publicly available at https://github.com/DrunkenPoet/GHAP
 Figure 1: Visual comparison. When reducing the number of Gaussians by 90%, our method outperforms other compaction techniques, such as PUP-3DGS, and remains competitive with the original 3DGS.
 1 Introduction
-Real-time 3D scene reconstruction and rendering dynamically generates photorealistic 3D representations from sensor data (e.g., multi-view images, LiDAR) with minimal latency, enabling critical applications in augmented/virtual reality (AR/VR), autonomous navigation, and immersive media [1, 2].
-The current state-of-the-art, 3D Gaussian Splatting (3DGS) [3], iteratively learns 3D anisotropic Gaussian primitives with color and opacity attributes to model these scenes.
+Real-time 3D scene reconstruction and rendering dynamically generates photorealistic 3D representations from sensor data (e.g., multi-view images, LiDAR) with minimal latency, enabling critical applications in augmented/virtual reality (AR/VR), autonomous navigation, and immersive media (fei20243d, ; chen2024survey, ).
+The current state-of-the-art, 3D Gaussian Splatting (3DGS) (kerbl20233d, ), iteratively learns 3D anisotropic Gaussian primitives with color and opacity attributes to model these scenes.
 During rendering, these 3D Gaussians are projected to 2D screens and $\alpha$-blended to achieve real-time photorealistic synthesis.
-However, 3DGS faces significant efficiency challenges: its iterative densification process often produces millions of redundant Gaussians for complex scenes [4, 5, 6].
-This inefficiency leads to high memory/storage costs and increased per-frame rendering time, limiting deployment on resource-constrained platforms like mobile and AR/VR devices [7].
+However, 3DGS faces significant efficiency challenges: its iterative densification process often produces millions of redundant Gaussians for complex scenes (lee2024compact, ; niemeyer2024radsplat, ; zhang2024lp, ).
+This inefficiency leads to high memory/storage costs and increased per-frame rendering time, limiting deployment on resource-constrained platforms like mobile and AR/VR devices (ren2024octree, ).
 Figure 2: Comparison of heuristic pruning and our method.
@@
 Our method better preserves the overall structure.
-A common strategy to improve efficiency is compaction [8, 9, 10, 7]—reducing the number of Gaussians while preserving rendering fidelity.
-With fewer primitives, both storage requirements and rendering efficiency (particularly per-frame rendering time) are naturally improved.
-This approach is viable because 3DGS densification inherently generates redundant primitives [4, 5, 6].
-Existing methods achieve compaction via pruning or random subset selection [11, 12, 5, 13, 14, 15, 6, 4, 16].
-These strategies, which are based on naïvely discarding information, are simple to implement, but may be ineffective.
-In particular, they tend to lose critical structural details (as shown in Figure 1) or distort the underlying geometry (as shown in Figure 2).
+A common strategy to improve efficiency is compaction (bagdasarian20243dgs, ; cheng2024gaussianpro, ; mallick2024taming, ; ren2024octree, )—reducing the number of Gaussians while preserving rendering fidelity.
+Fewer primitives result in reduced storage needs and faster rendering, improving per-frame performance.
+This approach is viable because 3DGS densification inherently generates redundant primitives (lee2024compact, ; niemeyer2024radsplat, ; zhang2024lp, ).
+Existing methods achieve compaction via pruning or random subset selection (fan2024lightgaussian, ; fang2024mini, ; niemeyer2024radsplat, ; liu2024atomgs, ; ali2024trimming, ; hanson2024pup, ; zhang2024lp, ; lee2024compact, ; chen2024hac, ).
+These strategies naïvely discard Gaussian primitives; while simple to implement, they are often ineffective.
+In particular, they tend to lose critical structural details (as shown in Fig.
+1) or distort the underlying geometry (as shown in Fig.
+2).
 Such losses can degrade rendering quality, particularly in regions with fine-scale features or complex material properties.
@@
 This two-stage strategy ensures that the compacted model maintains both geometric accuracy and photorealistic rendering quality.
-For GMR, we minimize the composite transportation divergence [17], which is rooted in optimal transport theory [18, 19] and allows an effective algorithm.
+For GMR, we minimize the composite transportation divergence (zhang2023gaussian, ), which is rooted in optimal transport theory (peyre2019computational, ; villani2009optimal, ) and allows an effective algorithm.
 We tailor this GMR algorithm (detailed in Section 3.2.1) so that it scales well in scenes like 3DGS with an extensive amount of Gaussians.
@@
   • We open a new pathway to view 3DGS representations as a Gaussian mixture and perform compaction from the perspective of Gaussian mixture reduction via optimal transport.
-  • We decouple geometry and appearance optimization in a scalable blockwise optimization scheme.
-  • Empirical results demonstrate that our method preserves rendering quality using only 10% Gaussians.
-  • We provide a plug-and-play module compatible with any 3DGS variant at any training stage.
+This contrasts with prior compaction methods that ignore geometric structure and often produce distortions, whereas our approach preserves geometric fidelity, offering a new and impactful direction for 3DGS.
+  • We are the first to adapt GMR to 3DGS.
+We introduce a novel cost function that yields closed-form, low-cost updates.
+We also develop a block-wise GMR algorithm guided by a KD-tree, enabling efficient large‑scale scene compaction.
+These strategies are non-trivial and bridges theory with practical scalability.
+  • Our method is post-hoc and compatible with any existing 3DGS pipeline, making it highly practical and broadly applicable.
+With minimal overhead, our approach achieves SOTA compaction performance, both in quality and efficiency.
+  • Empirical results demonstrate that our method preserves rendering quality at 10% retention ratio.
 2 Related Works
-Compaction Techniques in 3DGS 3DGS compaction seeks to minimize the Gaussian count while preserving image quality.
+Compaction Techniques in 3DGS.
+3DGS compaction seeks to minimize the Gaussian count while preserving image quality.
 Existing work falls into two operations: densification (where to add) and pruning (what to drop).
 Most current approaches rely on per-Gaussian heuristic scores.
-For densification, Taming-3DGS [10] ranks candidate Gaussians by combining gradient, pixel coverage, per-view saliency, and core attributes including opacity, depth, and scale; Color-cued GS [20] considers the view-independent spherical harmonics coefficient gradient to better capture color cues; and GaussianPro [9] guides growth using depth and normal maps.
-In the pruning phase, LightGaussian [11], Mini-Splatting [12], RadSplat [5], and AtomGS [13] compute an importance score from each Gaussian’s accumulated ray contribution, typically a mix of volume, opacity, transmittance, and hit count, and discard the lowest-ranked Gaussians.
-Gradient-aware variants prune by per-Gaussian gradients (Trimming-the-Fat) [14] or the second-order sensitivity score derived from the Hessian matrix (PUP-3DGS) [15].
-The score can also be trained via a learnable mask, as in LP-3DGS [6], Compact3DGS [4], and HAC [16].
-Moreover, multi-view consistency criteria discard Gaussians unseen by keyframes [21] or visible only in real but not virtual views [22].
-For a comprehensive survey of 3DGS compression and compaction, we refer readers to [8, 23].
+For densification, Taming‑3DGS (mallick2024taming, ) ranks candidate Gaussians by combining gradient, pixel coverage, per-view saliency, and core attributes including opacity, depth, and scale; Color-cued GS (kim2024color, ) considers the view-independent spherical harmonics coefficient gradient to better capture color cues; and GaussianPro (cheng2024gaussianpro, ) guides growth using depth and normal maps.
+In the pruning phase, LightGaussian (fan2024lightgaussian, ), Mini-Splatting (fang2024mini, ), RadSplat (niemeyer2024radsplat, ), and AtomGS (liu2024atomgs, ) compute an importance score from each Gaussian’s accumulated ray contribution, typically a mix of volume, opacity, transmittance, and hit count, and discard the lowest-ranked Gaussians.
+Gradient-aware variants prune by per-Gaussian gradients (Trimming-the-Fat) (ali2024trimming, ) or the second-order sensitivity score derived from the Hessian matrix (PUP-3DGS) (hanson2024pup, ).
+The score can also be trained via a learnable mask, as in LP-3DGS (zhang2024lp, ), Compact3DGS (lee2024compact, ), and HAC (chen2024hac, ).
+Moreover, multi-view consistency criteria discard Gaussians unseen by keyframes (matsuki2024gaussian, ) or visible only in real but not virtual views (ji2024neds, ).
+For a comprehensive survey of 3DGS compression and compaction, we refer readers to (bagdasarian20243dgs, ; bao20253d, ).
 Despite their success, most existing strategies evaluate each Gaussian independently, leaving open the question of whether the retained set is truly the best global surrogate.
 Our work addresses this gap from a probabilistic perspective via Gaussian mixture reduction.
-3DGS from Probabilistic Distribution Point of View Kheradmand et al.
-[24] formulate 3DGS as a Markov chain Monte Carlo process and use stochastic gradient Langevin dynamics to migrate dropped Gaussians onto retained ones, partially recycling lost information.
+3DGS from Probabilistic Distribution Point of View.
+Kheradmand et al.
+(kheradmand20243d, ) formulate 3DGS as a Markov chain Monte Carlo process and use stochastic gradient Langevin dynamics to migrate dropped Gaussians onto retained ones, partially recycling lost information.
 However, their update remains pairwise and lacks convergence guarantees under a principled divergence.
@@
 3 Method: Gaussian Herding across Pens
-Figure 3: An illustration of the proposed GHAP-3DGS approach.
+Figure 3: An illustration of the proposed GHAP approach.
 The process begins with full-resolution 3DGS training to obtain initial geometric and appearance features.
-These Gaussians are then spatially partitioned using a KD-tree and grouped into blocks—analogous to sheep pens.
-We then perform blockwise Gaussian Model Reduction (GMR) to generate a small number of representative Gaussians within each block.
-This step is analogous to the popular kernel herding method  [25].
+These Gaussians are then spatially partitioned using a KD-tree and grouped into blocks–analogous to sheep pens.
+We then perform blockwise Gaussian Mixture Reduction (GMR) to approximate the geometric shape within each block using a much smaller number of Gaussians.
+This step is analogous to the popular kernel herding method  (DBLP:journals/corr/abs-1203-3472, ).
 Finally, a lightweight appearance refinement step further optimizes the appearance feature of the reduced set.
-This multi-stage pipeline progressively guides the Gaussians in each block—analogous to herding across pens—toward a compact and high-fidelity representation.
-We present a principled approach for compacting 3DGS representations through Gaussian Mixture Reduction (GMR).
-Our key insight is that The 3DGS densification strategy, while effective for view synthesis, produces redundant Gaussian components that can be consolidated without sacrificing visual quality.
-We achieve this by interpreting the 3D scene surface as an unnormalized probability density function (PDF), where the trained 3DGS representation serves as a Gaussian mixture approximation of this PDF.
-Thus, our problem reduces to finding a Gaussian mixture with fewer components to approximate the original, redundant one.
-Our procedure separately optimizing geometric (position $\mu$, covariance $\Sigma$) and appearance (opacity $\alpha$, color $c$) attributes to ensure rendering fidelity.
-The key idea of our method is illustrated in Figure 3.
+This multi-stage pipeline progressively guides the Gaussians in each block–analogous to herding across pens–toward a compact and high-fidelity representation.
 3.1 Probabilistic Scene Representation
@@
 $\phi_{n}(x)=\sum_{i=1}^{n}\alpha_{i}\phi(x;\mu_{i},\Sigma_{i}),$
-[…]
+where $\alpha_{i}$ are the mixture weights satisfying $\alpha_{i}>0$ [^1] In statistics, Gaussian mixtures are defined with the constraint $\sum_{i=1}^{n}\alpha_{i}=1$ to ensure the area under the density function is $1$.
+However, we relax this constraint and consider unnormalized Gaussian mixtures, as in 3DGS, where the integral under the geometric surface need not equal $1$..
+In the context of 3DGS, let $x\in{\mathbb{R}}^{3}$ be spatial coordinates and $f(x)$ represent the implicit surface function (i.e., geometric shape and opacity that excludes color).
+The training process in 3DGS learns opacity parameter $\alpha_{i}$, location parameter $\mu_{i}$, and shape parameter $\Sigma_{i}$ such that
 $\phi_{n}(x)\approx f(x),\quad\forall x\in{\mathcal{X}},$
@@
 3.2 Compaction via Optimal Transport
-Motivated by the observation that many 3DGS algorithms [4, 5, 6] produce a significant number of redundant Gaussians during training, we improve rendering efficiency through compaction—reducing the number of Gaussian primitives to achieve lower memory usage and faster rendering while preserving visual fidelity.
+Motivated by the observation that many 3DGS algorithms (lee2024compact, ; niemeyer2024radsplat, ; zhang2024lp, ) produce a significant number of redundant Gaussians during training, we improve rendering efficiency through compaction—reducing the number of Gaussian primitives to achieve lower memory usage and faster rendering while preserving visual fidelity.
 The process consists of two key phases:
   1. 1.
-    Geometric Compaction via GMR: Leveraging our probabilistic interpretation, we formulate compaction as Gaussian Mixture Reduction (GMR) [26], approximating the original Gaussian mixture with redundant components with fewer components.
+    Geometric Compaction via GMR: Leveraging our probabilistic interpretation, we formulate compaction as Gaussian Mixture Reduction (GMR) (crouse2011look, ), approximating the original Gaussian mixture with redundant components by one with fewer components.
 This yields a compacted geometric representation: $\bar{\phi}_{m}(x)=\sum_{j=1}^{m}\bar{\alpha}_{j}\phi(x;\bar{\mu}_{j},\bar{\Sigma}_{j})$, where $m\ll n$.
@@
 3.2.1 Geometric Compaction via GMR
-Following Zhang et al. [17], we formulate compaction as minimizing the composite transportation divergence (CTD) between two Gaussian mixtures:
+Following Zhang et al. (zhang2023gaussian, ), we formulate compaction as minimizing the composite transportation divergence (CTD) between two Gaussian mixtures:
 Definition 1 (Composite transportation divergence).
@@
 ${{\mathcal{T}}}_{c}(\phi_{n},\phi^{\prime}_{m})=\inf\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}\pi_{ij}c(\phi(\cdot;\mu_{i},\Sigma_{i}),\phi(\cdot;\mu^{\prime}_{j},\Sigma^{\prime}_{j})):\sum_{j=1}^{m}\pi_{ij}=\alpha_{i},\sum_{i=1}^{n}\pi_{ij}=\alpha^{\prime}_{j}\right\}.$
-The CTD generalizes optimal transport [18] to mixtures, treating each component as a discrete distribution in Gaussian space.
+The CTD generalizes optimal transport (peyre2019computational, ) to mixtures, treating each component as a discrete distribution in the space of Gaussian distributions.
 The cost function measures the cost of moving one unit of Gaussian from one location to another, and $\pi_{ij}$ measures the corresponding amount of mass that is being moved.
@@
 With this formulation, the Gaussian mixture after compaction has optimal guarantee. The solution also guides the choice of $m$ to balance compactness and fidelity.
+As shown in Zhang et al. (zhang2023gaussian, ), (1) can be solved using the effective iterative algorithm in Algorithm 1.
 Algorithm 1 GMR via $k$-means Clustering
@@
 3:  Assignment Step:
-4:     for $i=1$ to $n$ do
+4:  for $i=1$ to $n$ do $\triangleright$ $\mathcal{O}(nm)$
 5:   Assign component $i$ to cluster ${\mathcal{C}}_{j}$ that minimizes $c(\phi(\cdot;\mu_{i},\Sigma_{i}),\phi(\cdot;\bar{\mu}_{j}^{(t-1)},\bar{\Sigma}_{j}^{(t-1)}))$
@@
 7:  Update Step:
-8:     for $j=1$ to $m$ do
+8:  for $j=1$ to $m$ do $\triangleright$ $\mathcal{O}(nm)$
 9:   Compute new cluster center: $\bar{\mu}_{j}^{(t)},\bar{\Sigma}_{j}^{(t)}=\operatorname*{arg\,min}\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}c(\phi(\cdot,\mu_{i},\Sigma_{i}),\phi(\cdot,\mu,\Sigma))$
@@
 11:  if no change in assignments then
-12:         for $j=1$ to $m$ do
+12:   for $j=1$ to $m$ do $\triangleright$ $\mathcal{O}(n)$
 13:     Compute weight: $\bar{\alpha}_{j}=\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}$
@@
 17:end for
-As shown in Zhang et al.
-[17], (1) can be solved using the effective iterative algorithm in Algorithm 1.
 The algorithm reduces to a $k$-means variant in Gaussian space: 1) The assignment step follows the same principle as traditional $k$-means, but replaces the $L^{2}$ distance between vectors with a cost function $c(\cdot,\cdot)$ between Gaussian distributions.
@@
 While the standard GMR algorithm provides optimal theoretical guarantees, its direct application to 3DGS compaction proves computationally prohibitive.
-Although the algorithm must converge in finite steps [17] [^2] In particular, we find that in our experiment, the algorithm converges in only around 6 iterations., the assignment step involves $nm$ evaluations of the cost function per iteration.
+Although the algorithm must converge in finite steps (zhang2023gaussian, ) [^2] In particular, we find that in our experiment, the algorithm converges in only around 6 iterations., the assignment step involves $nm$ evaluations of the cost function per iteration.
 In typical 3D scenes, the number of Gaussians scales as $n=\Omega(10^{5})$ [^3] The $a_{n}=\Omega(b_{n})$ ($O(b_{n})$) if there exist $C\geq 0$ such that $a_{n}\geq Cb_{n}$ ($a_{n}\leq Cb_{n}$)., and even after 95% reduction, each iteration would still require at least $10^{8}$ operations and memory storage.
@@
 The KL divergence considered in Zhang et al.
-[17] suffers from its key limitations: it requires computing inverse covariance matrices, which introduces (1) numerical instability due to small eigenvalues of covariance matrices, and (2) significant overhead despite operating in ${\mathbb{R}}^{3}$ (with $O(\rho s^{2}\log n)$ inversions needed).
+(zhang2023gaussian, ) suffers from (a) significant overhead of computing $O(\rho s^{2}\log n)$ covariance matrices inversions, and (b) numerical instability due to small eigenvalues of covariance matrices.
 To overcome this challenge, we introduce two key optimizations designed for 3DGS:
@@
 As demonstrated in Remark 1, this blockwise approach yields significant computational savings.
-While both KD-trees [27] and Octrees [28] are effective for spatial partitioning in 3D space, we employ a KD-tree for two key advantages.
+While both KD-trees (bentley1975multidimensional, ) and Octrees (meagher1982geometric, ) are effective for spatial partitioning in 3D space, we employ a KD-tree for two key advantages.
 First, it produces more balanced partitions across regions.
@@
 Furthermore, the reduction steps can be executed in parallel across blocks, offering additional computational speedup.
-  • Efficient Cost Function: We propose a novel cost function that is both computationally efficient and maintains approximation quality, overcoming limitations of KL divergence used in [17].
-We propose to use the following divergence:
+  • Efficient Cost Function: We introduce a novel cost function that overcomes the limitations of the KL divergence used in (zhang2023gaussian, ) by being computationally efficient without sacrificing approximation quality.
+Our proposed divergence is:
     $c(\phi(\cdot;\mu,\Sigma),\phi(\cdot;\mu^{\prime},\Sigma^{\prime}))=\|\mu-\mu^{\prime}\|_{2}^{2}+\|\Sigma-\Sigma^{\prime}\|_{F}^{2},$ (2)
-    which has three key advantages: First, it preserves distributional similarity as Gaussians are uniquely determined by their mean and covariance.
-Second, the assignment step requires only efficient vector/matrix norm computations.
-Third, the update step simplifies to weighted averages:
-    $\bar{\mu}_{j}^{(t)}=\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}\mu_{i}/\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i},\quad\quad\quad\bar{\Sigma}_{j}^{(t)}=\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}\Sigma_{i}/\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}.$
-3.2.2 Appearance Optimization
+    which offers three significant advantages: First, it preserves distributional similarity, as Gaussian distributions are uniquely determined by their mean and covariance.
+Second, the assignment step requires only efficient vector and matrix norm computations.
+Third, the update step simplifies to calculating weighted averages, thereby avoiding the computationally expensive covariance matrix inversions required by the KL divergence:
+    $\bar{\mu}_{j}^{(t)}=\frac{\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}\mu_{i}}{\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}},\quad\quad\quad\bar{\Sigma}_{j}^{(t)}=\frac{\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}\Sigma_{i}}{\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}}.$
 Algorithm 2 GHAP: 3DGS Compaction via Blockwise GMR
@@
 3:Stage 1: Geometric Compaction
-4:1. Build KD-tree from Gaussian centers $\{\mu_{i}\}_{i=1}^{n}$ with depth $d=\lfloor\log_{2}(n/s)\rfloor$
-5:2. For each leaf block ${\mathcal{B}}_{k}$:
+4:1. Build KD-tree from Gaussians $\{\mu_{i}\}_{i=1}^{n}$ with depth $d=\lfloor\log_{2}(n/s)\rfloor$ $\triangleright$ $\mathcal{O}(nd\log n)$
+5:2. For each leaf block ${\mathcal{B}}_{k}$: $\triangleright$ $\mathcal{O}(nmT/2^{d}$)
 6: Run Algorithm 1 to reduce to $\rho s$ Gaussians
@@
 8:1. Initialize appearance for each $\bar{\phi}_{j}$:
-9:$\bar{c}_{j}\leftarrow c_{i^{*}}$ and $\bar{\alpha}_{j}\leftarrow\alpha_{i^{*}}$ where $i^{*}=\operatorname*{arg\,min}_{i\in[n]}\|\mu_{i}-\bar{\mu}_{j}\|_{2}$
+[…]
 10:2. Fine-tune $\{\bar{\alpha}_{j},\bar{c}_{j}\}$ using standard 3DGS rendering pipeline for $T$ steps
-Following geometric compaction, we initialize the appearance attributes (opacity and colors) of the compacted Gaussian primitives.
+3.2.2 Appearance Optimization
+Following geometric compaction, we initialize the appearance attributes (opacity and color) of the compacted Gaussian primitives.
 For each primitive in the reduced mixture, we assign the appearance parameters from its closest counterpart in the original Gaussian mixture.
-Using these initial values, we then optimize the appearance attributes through backpropagation following the standard 3DGS training pipeline.
-We optimize the opacity instead of directly using the opacity from the GMR algorithm since these output weights not necessarily satisfy the constraint on the opacity that it must be within 00 and $1$.
-Finetuning the the opacity can achieve better visualization performance.
+Using these initial values, we then optimize the appearance attributes through backpropagation within the standard 3DGS training pipeline used in the first stage.
+We optimize the opacity instead of directly using the values from the GMR algorithm because its output weights do not necessarily satisfy the constraint that opacity must be between 0 and $1$.
+Fine-tuning the opacity leads to better visualization performance.
 3.3 Training Details with GHAP Algorithm
@@
 The process begins with standard 3DGS optimization for $T$ steps, followed by blockwise GMR.
-We then freeze the geometric parameters ($\mu$ and $\Sigma$) while fine-tuning appearance attributes (opacity $\alpha$ and color $c$) through an additional $T$-step 3DGS optimization.
-This procedure can be iteratively applied throughout training as needed.
+We then freeze the geometric parameters ($\mu$ and $\Sigma$) while fine-tuning the appearance attributes (opacity $\alpha$ and color $c$) through an additional $T$-step 3DGS optimization.
+This procedure can be applied iteratively throughout training as needed.
 4 Experiments
 4.1 Experimental Setup
-Datasets
-For a comprehensive evaluation of our GHAP algorithm, we use three real-world datasets: Tanks and Temples [29], Mip-NeRF 360 [30], and Deep Blending [31], which cover varying levels of detail, lighting conditions, and scene complexities.
-For each dataset, we adopt the same scenes as in [8].
-  • Tanks and Temples: We evaluate two unbounded outdoor scenes, “Truck” and “Train”, both featuring centered viewpoints.
+Datasets.
+For a comprehensive evaluation of GHAP algorithm, we use three real-world datasets: Tanks & Temples (knapitsch2017Tanks, ), Mip-NeRF 360 (hedman2018deep, ), and Deep Blending (barron2022mip, ), which cover varying levels of detail, lighting conditions, and scene complexities.
+For each dataset, we adopt the same scenes as in (bagdasarian20243dgs, ).
+  • Tanks & Temples: We evaluate two unbounded outdoor scenes, “Truck” and “Train”, both featuring centered viewpoints.
   • Mip-NeRF 360: We test on a mix of indoor and outdoor scenes, including “Bicycle”, “Bonsai”, “Counter”, “Flowers”, “Garden”, “Kitchen”, “Room”, “Stump”, and “Treehill”, all with centered viewpoints.
   • Deep Blending: We include two indoor scenes, “Dr. Johnson” and “Playroom”, where the viewpoint is directed outward.
-Baselines
-To evaluate the effectiveness of our proposed method, we compare it against several strong baselines in the domain of 3DGS compaction: Mini-Splatting-D [12], AtomGS [13], LightGaussian [11], and PUP-3DGS [15].
-We set $s=3\times 10^{3}$.
-The sensitivity analysis of $s$ is given in Appendix.
-The initialization of the iterative algorithm in (1) randomly selects $m$ Gaussians per block.
-Evaluation Metrics We assess 3DGS compaction using standard metrics for rendering quality and efficiency.
-For quality, we report: (1) PSNR, measuring pixel-level accuracy; (2) SSIM, evaluating perceptual similarity based on luminance, contrast, and structure; and (3) LPIPS, capturing perceptual distance via a learned model.
+Baselines.
+To evaluate the effectiveness of our proposed method, we compared it against four strong compaction techniques: LightGaussian (fan2024lightgaussian, ), PUP-3DGS (hanson2024pup, ), Trimming the Fat (ali2024trimming, ), and MesonGS (xie2024mesongs, ), as well as four end-to-end 3DGS variants: Mini-Splatting(-D) (fang2024mini, ), AtomGS (liu2024atomgs, ), 3DGS-MCMC (kheradmand20243d, ), and LocoGS (shin2025locality, ).
+Notably, Mini-Splatting-D and AtomGS were employed as backbone models in our approach, while the other variants were used for direct comparative evaluation against the compaction methods.
+A consistent evaluation protocol was established to ensure fair and reliable conclusions.
+For post-training compaction methods applicable to pre-trained models–including GHAP, LightGaussian, PUP-3DGS, Trimming the Fat, and MesonGS–we initialized all from the same backbone model (trained using vanilla 3DGS, Mini-Splatting-D, or AtomGS for 15k iterations) and applied their respective compaction procedures directly, excluding any compression-specific modules.
+All models subsequently underwent identical fine-tuning for 15k iterations to achieve the target retention ratio.
+For the other end-to-end variants (MiniSplatting, 3DGS-MCMC and LocoGS), we executed training for 30k iterations under their default configurations.
+Detailed experimental steps for each method can be found in the appendix.
+Evaluation Metrics.
+We assess 3DGS compaction using standard metrics for rendering quality.
+We report: (1) PSNR, measuring pixel-level accuracy; (2) SSIM, evaluating perceptual similarity based on luminance, contrast, and structure; and (3) LPIPS, capturing perceptual distance via a learned model.
 Higher PSNR/SSIM and lower LPIPS indicate better quality.
-For efficiency, we measure FPS across retention ratios, where lower retention yielding higher FPS implies greater rendering efficiency.
-Device
-All experiments were conducted on a server with 256 GB RAM and a 96-core Intel® Xeon® Platinum 8255C CPU, and on a workstation equipped with five NVIDIA RTX 3090 GPUs, each with 24 GB of VRAM.
+For each method, we also report the corresponding number of Gaussian primitives.
 4.2 Quantitative Results
-We first demonstrate the effectiveness of our GHAP compaction method as a plug-in module within 3DGS and its variants.
-Second, we show that our approach outperforms other SOTA compaction techniques.
-We present the results first and summarize our findings at the end of the section.
-As a Plug-In Compaction Method
+We first show that our approach outperforms other SOTA compaction techniques.
+Second, we demonstrate the effectiveness of our GHAP compaction method as a plug-in module within 3DGS and its variants.
+We present the experimental results followed by our key findings.
+Comparison with SOTA. We compare our method against a comprehensive set of baselines, which can be categorized into two groups:
+Our approach is evaluated in two configurations: 3DGS+GHAP and MiniSplatting+GHAP.
+The former uses the same vanilla 3DGS backbone and 10% retention rate as the pruning baselines for a direct comparison of compaction strategies.
+The latter replaces the built-in pruning step in Mini-Splatting with our GHAP algorithm to demonstrate its effectiveness on a different backbone.
+Quantitative results are summarized in Table LABEL:tab:sota_comparison, with methods grouped by their final primitive count.
+As expected, LocoGS achieves strong performance due to its larger primitive count.
+Among methods with comparable primitive counts (Group 2), our 3DGS+GHAP achieves superior performance in SSIM and PSNR, with a marginally lower LPIPS score.
+The advantage of our compaction strategy is further evident when using the Mini-Splatting backbone.
+Our MiniSplatting+GHAP outperforms other compaction-based approaches while often using fewer primitives.
+As shown in Fig.
+4 (left), this performance lead is consistent across a wide range of retention ratios, not just at $\rho=0.1$.
+Table 1: Quantitative comparison of rendering quality against SOTA methods.
+Our method (GHAP), applied to two different backbones (3DGS and Mini-Splatting), consistently matches or surpasses pruning-based baselines, even at a low retention rate (10%), while maintaining competitive runtime and memory usage.
+The best results for a given number of Gaussians are shown in bold; second best are underlined.
+Method Tanks&Temples MipNeRF-360 Deep Blending
+[…]
+Vanilla 3DGS 0.853 23.785 0.169 1577 0.813 27.554 0.221 2627 0.907 29.816 0.238 2475
+LocoGS 0.843 23.655 0.191 571 0.798 27.049 0.257 674 0.903 29.972 0.261 529
+3DGS+GHAP (ours) 0.818 23.312 0.242 157 0.764 26.404 0.314 263 0.905 29.647 0.264 248
+LightGaussian 0.756 22.113 0.306 158 0.735 25.674 0.331 263 0.869 28.010 0.327 248
+PUP-3DGS 0.767 21.519 0.280 158 0.753 25.332 0.309 262 0.895 29.153 0.274 248
+Trimming the Fat 0.776 21.535 0.293 156 0.731 25.255 0.343 263 0.887 28.056 0.302 247
+MesonGS 0.811 20.714 0.208 157 0.773 24.924 0.264 263 0.896 28.693 0.264 248
+3DGS-MCMC 0.779 22.141 0.282 157 0.763 25.957 0.309 263 0.885 28.976 0.298 248
+MiniSplatting 0.799 22.661 0.265 78 0.759 26.022 0.318 111 0.895 29.395 0.289 125
+MiniSplatting+GHAP (ours) 0.835 23.232 0.198 79 0.802 27.090 0.250 112 0.909 30.042 0.254 127
+Figure 4: Comparison of compaction methods on the Tanks & Temples dataset. Left: Rate-Distortion (RD) curves; middle: computational time, and right: memory consumption.
+Runtime & Memory Usage Comparison.
+Crucially, the improved performance of our method does not come at a computational cost.
+As depicted in Fig.
+4 (middle), our method’s runtime is faster than all baselines except the exceptionally swift Trimming the Fat.
+While our method exhibits a slightly higher memory footprint (in Fig.
+4 right) during compaction due to pairwise distance computation in each KD-tree block, the difference is not substantial (less than an order of magnitude).
+As a Plug-In Compaction Method.
 Our method can be used as a plug-in compaction method in various 3DGS training algorithms.
@@
 To verify this, we apply our compaction method within various 3DGS pipelines.
-Specifically, we consider three representative variants: 3DGS-30K, AtomGS, and Mini-Splatting-D.
+Specifically, we consider three representative variants as the backbone: 3DGS (kerbl20233d, ) , AtomGS (liu2024atomgs, ), and Mini-Splatting-D (fang2024mini, ).
 Each of them employs a distinct densification strategy, and our method can be directly embedded into the pipeline without extensive engineering effort.
-For each variant, we evaluate performance under two retention ratios (10% and 20%) and report results using three standard metrics: PSNR, SSIM, and LPIPS. Table 1 summarizes our experimental results.
-Table 1: Quantitative evaluation of our GHAP method with respect to compaction performance applied to different 3DGS variants under varying retention ratios (10% and 20%).
-Method $\rho$ Tanks&Temples MipNeRF-360 Deep Blendering
+For each backbone, we evaluate performance under two retention ratios (10% and 20%).
+Tab.
+2 summarizes our experimental results.
+Table 2: Quantitative results with different backbones. The compaction performance of our GHAP method when used with different 3DGS variants as backbones under varying retention ratios.
+Backbone $\rho$ Tanks&Temples MipNeRF-360 Deep Blending
 […]
-3DGS-30k 10% 0.818 23.312 0.242 0.764 26.404 0.314 0.905 29.647 0.264
-20% 0.835 23.615 0.212 0.788 26.973 0.275 0.907 29.864 0.252
-100% 0.853 23.785 0.169 0.813 27.554 0.221 0.907 29.816 0.238
-Mini-Splatting-D 10% 0.835 23.232 0.198 0.802 27.090 0.250 0.909 30.042 0.254
-20% 0.855 23.403 0.171 0.821 27.310 0.214 0.912 30.170 0.238
-100% 0.848 23.338 0.140 0.832 27.486 0.176 0.907 29.980 0.211
-AtomGS 10% 0.793 22.988 0.274 0.764 26.535 0.307 0.899 29.347 0.282
-20% 0.812 23.282 0.240 0.788 27.025 0.269 0.902 29.347 0.268
-100% 0.814 23.289 0.235 0.796 27.135 0.251 0.902 29.314 0.267
-Comparison with SOTA Compaction Methods
-We compare our method with two recent compaction baselines: PUP-3DGS and LightGaussian.
-To ensure a fair evaluation, all methods are built upon the same backbone model, 3DGS-30K.
-Compaction is performed at the same stage, when the number of Gaussians reaches 15K, followed by 15K fine-tuning iterations.
-During fine-tuning, we use identical optimization settings across all methods, including learning rate and other hyperparameters.
-Two compaction ratios are considered: 10% and 20%.
-A detailed comparison of rendering quality across different compaction methods and retention ratios is presented in Table 2.
-Table 2: We compare our GHAP method against LightGaussian and PUP-3DGS, modifying their original workflows to align with our compaction pipeline.
-Method-Compact Rate Tandt&Temples MipNeRF-360 Deep Blendering
-[…]
-3DGS-30k-100% 0.853 23.785 0.169 0.813 27.554 0.221 0.907 29.816 0.238
-GHAP(Ours)-10% 0.818 23.312 0.242 0.764 26.404 0.314 0.905 29.647 0.264
-Light-GS-10% 0.791 22.560 0.248 0.779 26.585 0.261 0.861 27.136 0.326
-Pup-GS-10% 0.767 21.519 0.28 0.753 25.332 0.309 0.895 29.153 0.274
-GHAP(Ours)-20% 0.835 23.615 0.212 0.788 26.973 0.275 0.907 29.864 0.252
-Light-GS-20% 0.791 22.559 0.248 0.779 26.588 0.261 0.861 27.142 0.326
-Pup-GS-20% 0.809 22.603 0.228 0.790 26.671 0.257 0.905 29.719 0.248
-Figure 4: Visualization before and after compaction. Various scenes before and after compaction under different 3DGS algorithms. Our GHAP method yields negligible loss in rendering quality.
-Takeaways
-Based on these experiments, we draw four key observations:
-  • Superior fidelity retention.
-Our method substantially outperforms both LightGaussian and PUP-3DGS in SSIM and PSNR.
-Surprisingly, our method can even achieve better performance than the original method, this is demonstrated in Table 1 on the Deep Blending scene with the Mini-Splatting-D backbone and at a 10% retention ratio.
-Likewise, in Table 2, we consistently surpass LightGaussian and PUP-3DGS in most datasets and metrics.
-  • Efficiency gains.
-At a 10% Gaussian primitives retention rate, our compaction framework yields a 140.3% increase in rendering speed (FPS), demonstrating that significant model slimming can be achieved without sacrificing and even often improving performance (see ablation study).
-  • Cross-dataset consistency.
-Performance remains stable across datasets.
-In outdoor scenes (Tanks and Temples), SSIM value fidelity drops ratio (relative drop in rendering metrics after compaction) less than 4.1%.
-Meanwhile, for the indoor scene, the PSNR value retains 99.4% of the original value.
-  • LPIPS sensitivity.
-We observe that our GMR compaction method incurs a relatively larger degradation in LPIPS compared to SSIM and PSNR, although the reduction remains within acceptable bounds.
-This trend aligns with prior findings [11, 15], as LPIPS is more sensitive to structural and textural distortions.
-Since compaction inherently reduces model capacity, fine-grained texture reconstruction becomes more challenging, which explains the behavior observed in our qualitative and quantitative analyses
-Qualitative Visualization
-We visualize two scenes for each baseline 3DGS algorithm before and after compaction in Figure 4.
-Across most scenes, our compaction method successfully preserves rendering quality using only 10% of the Gaussian primitives.
-By design, our reduced Gaussians approximate the original Gaussians from the 3DGS algorithm.
-If the original model lacks good visual quality due to lack to Gaussian primitives in certain regions, our method cannot perform well—as demonstrated in the ‘Bicycle’ scene, where the poor performance of the Atom-GS baseline prevents our approach from maintaining high fidelity after compaction.
-Interestingly, in some cases, our post-compaction results surpass the pre-compaction quality.
-A representative example is the ‘Kitchen’ scene based on Mini-Splatting-D.
-We conjecture this occurs because Mini-Splatting-D lacks a pruning mechanism, often generating an oversaturated set of Gaussians that introduce visual artifacts (e.g., unnatural shadows in the lower-left region).
-Our method mitigates this issue by globally regularizing the Gaussian distribution: improving expressiveness, reducing unnecessary density, and preserving the underlying 3D structure.
+3DGS-30k 10% 0.818 23.312 0.242 157 0.764 26.404 0.314 263 0.905 29.647 0.264 248
+20% 0.835 23.615 0.212 314 0.788 26.973 0.275 526 0.907 29.864 0.252 596
+100% 0.853 23.785 0.169 1577 0.813 27.554 0.221 2627 0.907 29.816 0.238 2475
+Mini-Splatting-D 10% 0.835 23.232 0.198 313 0.802 27.090 0.250 357 0.909 30.042 0.254 331
+20% 0.855 23.403 0.171 626 0.821 27.310 0.214 614 0.912 30.170 0.238 662
+100% 0.848 23.338 0.140 3132 0.832 27.486 0.176 3578 0.907 29.980 0.211 3316
+AtomGS 10% 0.793 22.988 0.274 189 0.764 26.535 0.307 293 0.899 29.347 0.282 271
+20% 0.812 23.282 0.240 378 0.788 27.025 0.269 586 0.902 29.347 0.268 542
+100% 0.814 23.289 0.235 1897 0.796 27.135 0.251 2928 0.902 29.314 0.267 2709
+The quantitative results in Tab.
+2 demonstrate that our method effectively preserves the backbone models’ visual quality, even at an extreme retention rate of 10%.
+Notably, on some scenes (highlighted in bold), the compacted model’s performance surpasses that of the uncompacted backbone.
+Beyond quantitative metrics, we provide a qualitative analysis by visualizing multiple scenes before and after compaction in Fig. 5.
+Figure 5: Visual Quality Before and After Compaction.
+Visual results for various scenes under different 3DGS backbones, compacted to 10% of their primitives using our GHAP method.
+Our approach preserves rendering quality with negligible loss.
+In some cases (e.g., “Kitchen”), compaction even improves quality by regularizing an over saturated Gaussian distribution.
+As evidenced in the figure, our method successfully preserves rendering quality across most scenes while using only 10% of the Gaussian primitives.
+Interestingly, in certain cases, our compaction not only preserves but surpasses the original quality.
+A representative example is the “Kitchen” scene (Mini-Splatting-D backbone).
+We conjecture this improvement occurs because Mini-Splatting-D lacks a pruning mechanism, often generating an over saturated set of Gaussians that introduce visual artifacts (e.g., the unnatural shadows in the lower-left region).
+Our method acts as a global regularizer, mitigating this issue by reducing unnecessary density while improving the overall expressiveness and preserving the underlying 3D structure.
+Naturally, our approach is inherently limited by the quality of its input.
+If the original model suffers from significant artifacts due to a lack of primitives in certain regions, our compaction cannot resolve these fundamental issues.
+This limitation is demonstrated in the “Bicycle” scene, where artifacts present in the Atom-GS backbone persist after compaction.
 4.3 Ablation Studies
-GMR versus Random Subsampling We evaluate the contribution of each stage in our pipeline—geometric compaction and appearance optimization—on the Tanks and Temples dataset.
-For comparison, we use random subsampling at the same compaction ratio.
-This allows us to assess the effectiveness of our design choices at each stage.
-The experiments show that our compaction method is significantly more effective than random subsampling,
+Influence of KD Tree Depth.
+To validate the effectiveness of the KD-tree partitioning strategy, we provide an ablation study on it.
+Due to the large scale of the three primary datasets, low KD-tree depths result in an excessive number of points per block, making it infeasible to run the GMR algorithm.
+Therefore, we conduct ablation experiments on the smaller mic scene from NeRF-Synthetic.
+Results in Fig.
+6 show that increasing KD-tree depth reduces memory usage and runtime, while PSNR first improves and then declines.
+This indicates that moderately finer partitions allow GMR to compact regions more effectively, whereas overly fine splits may fragment primitives and degrade quality.
+Figure 6: Hyperparameters. The impact of KD tree depth on runtime performance, memory footprint, and rendering quality.
+Loss Function Design.
+As shown in Tables LABEL:tab:sota_comparison and 2, our method exhibits a slight underperformance on the LPIPS metric.
+To address this, we investigate whether incorporating an LPIPS loss term can yield improvements.
+Our baseline loss function follows the vanilla 3DGS formulation, using an L1-to-SSIM ratio of 8:2.
+We experiment with two new weighting schemes that include LPIPS: (1) L1:SSIM:LPIPS = 8:1:1, and (2) L1:SSIM:LPIPS = 6:2:2.
+The results of this ablation study on the Tanks&Temples dataset are shown in Fig.
+7 (with additional results in the Appendix).
+Our analysis reveals a trade-off between perceptual and distortion metrics: increasing the weight of the LPIPS loss improves LPIPS scores but leads to a slight deduction in PSNR and SSIM.
+Figure 7: Comparison of different loss weights. A trade-off between perceptual and distortion metrics.
+GMR versus Random Subsampling.
+We conduct an ablation study on the Tanks&Temples dataset to evaluate the contribution of each stage in our pipeline: geometric compaction and appearance optimization.
+A random subsampling baseline is included to assess the effectiveness of our design choices.
 Table 3: Ablation study.
@@
 +15K Fine-tuning 0.712 21.312 0.282
-producing a more meaningful scene representation.
-Appearance optimization further leads to substantial quantitative improvements.
-Together, these results validate the necessity and effectiveness of both stages in our proposed pipeline.
-Figure 5: GHAP efficiency-quality trade-off. SSIM and FPS (top) and LPIPS and PSNR (bottom) across different retention ratios $\rho$.
-Rendering Metrics under GHAP
-To analyze the efficiency-quality trade-off of the proposed GHAP approach, we vary the retention ratio $\rho$ from $10\%$ to $100\%$.
-A lower retention ratio enables more aggressive compaction, resulting in faster rendering speed but potentially compromising rendering quality.
-Our results demonstrate that decreasing the retention ratio yields 2×2$\times$2 × faster rendering with only marginal quality degradation.
-For instance, as $\rho$ decreases from $100\%$ to $10\%$: LPIPS increases modestly from 0.14 to 0.21; SSIM declines slightly from 0.885 to 0.860; PSNR remains stable between 25.0 dB and 25.5 dB.
-These findings indicate that retaining even a modest number of Gaussians effectively preserves visual quality, achieving an excellent performance-fidelity trade-off.
-5 Conclusion
-We present a novel optimal transport-based global Gaussian reduction approach for 3D Gaussian Splatting, addressing the inefficiencies of existing Gaussian pruning and compaction methods.
-By formulating the problem as Gaussian mixture reduction, our method minimizes composite transport divergence to derive a compact geometric representation, followed by appearance fine-tuning with significantly fewer primitives.
-It provides a theoretically sound foundation for efficient neural rendering.
-Extensive experiments demonstrate that our approach achieves rendering quality nearly indistinguishable from vanilla 3DGS while retaining only  10% of the Gaussians, outperforming existing compaction techniques.
-Numerical results show our framework is versatile, compatible with various 3DGS pipelines.
-This work advances the practical deployment of 3DGS by striking an optimal balance between computational efficiency and visual fidelity.
-Future work includes extending our method to dynamic 3DGS for real-time temporal scenes and developing adaptive reduction strategies for task-specific needs.
+The results, presented in Table 3, compare two compaction schemes followed by the same fine-tuning procedure.
+Our findings demonstrate that: first, our compaction procedure is significantly more effective than random subsampling and other pruning baselines (as shown in Tab.
+LABEL:tab:sota_comparison).
+Second, the subsequent appearance optimization stage provides substantial quantitative improvements for both compaction approaches.
+The significant performance gain over the baseline validates the necessity and effectiveness of both stages in our proposed pipeline.
+5 Conclusion and Discussion
+We propose an optimal transport-based Gaussian mixture reduction framework for 3D Gaussian Splatting, achieving compact yet faithful representations.
+By minimizing composite transport divergence with appearance fine-tuning, our method preserves high visual fidelity while retaining only  10% of Gaussians, outperforming prior compaction techniques.
+The framework scales efficiently via block-wise KD-tree partitioning and integrates seamlessly with diverse 3DGS pipelines.
+Future directions include enhancing robustness across challenging scene types, incorporating perceptual objectives, developing multi-scale and overlap-aware partitioning, adopting auto-tuned schedules, and extending to dynamic 3DGS for real-time temporal rendering.
 Acknowledgments and Disclosure of Funding
-Qiong Zhang is supported by the National Natural Science Foundation of China Grant 12301391.
+Qiong Zhang is supported by the National Natural Science Foundation of China Grant 12301391. The authors would like to thank the anonymous reviewers for their constructive suggestions.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: Our main claim are summarized in Section 1 and Figure 1, Section 3 and Section 4 offer detailed explainations.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: We include the limitations of our work in Appendix.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: We have stated all assumptions explicitly within the theorem and include the complete proof of our theoretical results in Appendix.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: We explained our settings and hyperparameters in Section 4 and Appendix .
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: The data is open-source and the GitHub link is provided.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: We placed the detailed experimental settings of the experiments in Section 4 and the Appendix of the paper and the complete details with the code in the supplementary materials.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: We show results in Appendix.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: All experiments were conducted on a server with 256 GB RAM and 64 cores Intel ® Xeon ® Gold 6330 CPU, and a standard workstation equipped with a single NVIDIA RTX 3090 GPU.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: The research conducted in this paper complies with the NeurIPS Code of Ethics in all aspects.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49.
+Justification: This paper proposes a novel method of compaction for 3D Gaussian Splatting, which is broadly applicable to 3D real-time radiance field rendering.
+Its potential impact is further analyzed in Section 5.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: This paper involves no such risks.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We have ensured that all assets used in this paper are properly cited and their owners credited.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: Our implementation code is available in supplementary materials.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: This paper does not involve crowdsourcing nor human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: This paper does not involve crowdsourcing nor human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: The development of the core method in this research did not involve LLMs as any important, original, or non-standard components.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
 Appendix
@@
     An MM-based exhaustive algorithm with $O(m^{n})$ complexity solves (1).
-Appendix B More Background and Involved Techniques in Paper
-B.1 Gaussian splatting and efficient variants
-Due to its advantages in rendering speed and visual fidelity, 3DGS has been adopted in a broad range of applications, including human reconstruction, AI-generated content, autonomous driving, and beyond [34, 35, 36, 37].
-Extensions to dynamic 3DGS, editable 3DGS, and surface representation have further broadened its utility [38, 39, 40].
-However, the standard 3DGS suffers from significant storage, memory, and rendering costs when applied to large-scale or complex scenes.
-This limits its deployment on resource-constrained platforms and in broader areas.
-To improve storage and rendering efficiency while maintaining visual quality, recent research has focused on two complementary directions: compression and compaction [8].
-Compression methods shrink file size by exploiting redundancy in Gaussian attributes (e.g., using vector quantization to cluster similar parameters) [11, 41] or structuring Gaussians into more compact forms (e.g., grids, anchors, and shared codes) [42, 43].
-On the other hand, compaction aims to reduce the number of Gaussians required for accurate rendering by optimizing their spatial distribution and pruning redundant primitives [12, 6, 5].
-Hybrid pipelines such as LightGaussian [11] and Octree-GS [7] combine compression and compaction together to deliver highly compact and lightweight neural renders.
-B.2 Detailed 3DGS pipline
+Appendix B Experiment Steps and Complexity Analysis in More Detail
+B.1 Detailed 3DGS Pipline
 The advantages of 3DGS in rendering speed and image fidelity have made it applicable to a wide range of tasks, including human reconstruction, AI-generated content, autonomous driving, and beyond [36, 37, 38, 39].
@@
 During training, Gaussians are cloned (under-reconstruction) or split (over-reconstruction) based on view-space gradient magnitude, and low-contribution Gaussians are pruned.
-B.3 Involved 3DGS Variants
-Although 3DGS features a well-designed pipeline and an efficient differentiable rasterization scheme, its densification process still suffers from several limitations at the detail level.
-For example, the non-uniform distribution of Gaussians can lead to overlapping primitives in dense regions and under-reconstruction in sparse areas.
-In addition, the optimization tends to favor large Gaussians that quickly cover the scene to achieve fast convergence.
-This behavior, however, often results in poor reconstruction of high-frequency details and introduces geometric noise.
-Consequently, a growing body of work has focused on refining the densification stage of the original 3DGS framework.
-In this section, we highlight two such methods—Atom-GS [13] and Mini-Splatting (-D) [12]—which specifically enhance the densification process.
-While these methods also modify other components such as pruning, we isolate and utilize only the early-stage point cloud generated before 15K training iterations.
-Applying our compaction method to these two distinct variants further illustrates its generality.
-It is important to note that our approach is broadly applicable to any point cloud produced by 3DGS-style pipelines, as long as each Gaussian retains its mean, covariance, and opacity, along with a reasonable optimization strategy for opacity and color.
-Atom-GS
-To address the tendency of 3DGS to rely on large Gaussians for rapid scene coverage, AtomGS replaces the free-form Gaussian optimization in 3DGS with a structured formulation based on atom Gaussians—uniform, isotropic primitives that better adhere to the underlying scene geometry.
-In addition, AtomGS introduces geometry-guided optimization to further enhance the spatial fidelity of the reconstructed scene.
-AtomGS builds upon two key techniques:
-  1. 1.
-    Atomized Proliferation.
-    • Constrains Gaussians in high-detail regions into fixed-size, isotropic Atom Gaussians.
-    • Uses a geometric schedule to refine scale and enhance surface coverage over training iterations.
-    • Merges Gaussians in flat areas to reduce redundancy.
-  2. 2.
-    Geometry-Guided Optimization.
-    • Introduces Edge-Aware Normal Loss by weighting normal smoothness using an edge map derived from the RGB image.
-    • Preserves sharp features while smoothing flat regions, improving alignment with the true 3D geometry.
-Mini-Splatting (-D)
-The original 3DGS often suffers from non-uniform Gaussian distributions, resulting in excessive overlap in dense regions and insufficient reconstruction in sparse areas.
-Mini-Splatting addresses this issue by reorganizing the spatial distribution of Gaussians through a two-stage pipeline consisting of densification and simplification.
-Its goal is to reduce the total number of Gaussians without compromising rendering quality.
-Mini-Splatting-D is a variant of Mini-Splatting, tailored for quality-oriented rendering.
-While it follows a similar overall pipeline, the key difference lies in the omission of the simplification stage, allowing the method to focus entirely on maximizing rendering quality.
-In our implementation, we focus solely on the densification process used in our pipeline.
-  • Blur Split: Identifies Gaussians that dominate blurry regions (via screen-space analysis) and splits them to recover fine detail.
-  • Depth Reinitialization: Reprojects depth maps derived from midpoints between ray-ellipsoid intersections to generate dense and uniform 3D points for reinitialization.
-Beyond evaluating compatibility with different 3DGS densification variants, we compare our method against recent state-of-the-art compaction techniques.
-Specifically, we consider two representative baselines: LightGaussian [11] and PUP-3DGS [15].
-A brief overview of each method is provided below.
-LightGaussian
-Original 3DGS overpopulates Gaussians during densification, causing storage bloat and rendering inefficiency.
-LightGaussian prunes redundant Gaussians based on a global significance score (computed via hit count, opacity, and normalized volume), followed by fine-tuning to recover lost details.
-The steps include:
-  1. 1.
-    Compute significance for each Gaussian by a score based on the volume of the Gaussian sphere and the number of covered pixels.
-  2. 2.
-    Prune low-score Gaussians.
-  3. 3.
-    Fine-tune retained Gaussians with photometric loss.
-PUP-3DGS
-LightGaussian employs a visibility-aware importance score, which tends to favor pruning small Gaussians while retaining larger background primitives.
-In contrast, PUP-3DGS introduces a mathematically principled importance measure based on second-order derivatives, allowing it to preserve small but geometrically critical Gaussians in foreground regions.
-The technical steps are:
-  1. 1.
-    Fisher Approximation. Computes the gradients of the Gaussian parameters to approximate the Hessian matrix, which is then used to derive the sensitivity score.
-  2. 2.
-    Iterative Pruning. Rank Gaussians by sensitivity scores, prune low-sensitivity ones, and fine-tune iteratively to recover lost fidelity.
-Appendix C Sensitivity Analysis
-C.1 Choice of KD-tree Depth
-In our algorithm, the KD-tree depth $d$ (or equivalently, the number $s$ of Gaussians per KD block) is treated as a hyperparameter.
-We investigate how rendering quality varies with different values of $d$ on the Truck scene.
-The results are shown in Figure 6.
-Figure 6: Sensitivity Analysis of $d$. Our GHAP algorithm exhibits stable rendering performance across varying KD-tree depths $d$.
-We observe that, within a reasonable range, the KD-tree depth $d$ has minimal impact on rendering quality.
-In our experiments on the Truck scene, setting $d<6$ results in distance matrices that are too large to fit into GPU memory, causing memory overflow.
-Conversely, when $d>16$, the number of partitions becomes excessively large, and most blocks contain one or zero Gaussians, rendering the algorithm inapplicable.
-C.2 Sampling Iteration
-Since most 3DGS-based algorithms complete the densification stage before 15K training iterations, we apply our sampling procedure after this point to achieve a target retention ratio.
-To study the effect of sampling at different post-densification stages, we evaluate our method at various iterations beyond 15K.
-The experimental results are presented in Figure 7.
-Figure 7: Sensitivity Analysis of Sampling Iteration. The rendering performance of our GHAP algorithm has negligible loss by variations in sampling iteration.
-As shown in the results, sampling at any iteration beyond 15K has minimal impact on the final rendering quality.
-These results indicate that continuing optimization beyond 15K provides diminishing returns.
-Moreover, once sampling is performed, only a small number of fine-tuning iterations are needed for the compacted model to converge.
-Based on these observations, we recommend applying sampling uniformly at the 15K iteration, followed by an additional 15K fine-tuning iterations for all experiments.
-Appendix D Additional Scene Visualizations
-We report a more comprehensive set of results for the plug-in experiments in Table 4, including two additional metrics: the number of Gaussians (in thousands) and rendering speed measured in FPS.
-As shown in the updated results, our compaction method significantly improves FPS compared to the original models, often yielding over a 2$\times$ speedup in rendering performance.
+B.2 Algorithm Complexity
+For Algorithm 1, we discuss its computational cost for reducing $n$ Gaussians to $m$ Gaussians.
+The assignment step requires computing pairwise distances between all $n$ input Gaussians and $m$ cluster centers, resulting in a complexity of $\mathcal{O}(nm)$.
+The update step involves computing the barycenters.
+Since the assignments are already known and the update for each cluster is linear in the number of assigned Gaussians, this step has a complexity of $\mathcal{O}(n)$.
+The overall time complexity of Algorithm 1 is ${\mathcal{O}}(nm)$.
+Our GHAP algorithm consists of two components: Geometric Compaction and Appearance Optimization (fine-tuning).
+In this analysis, we focus only on the time complexity of the geometric compaction stage.
+The KD-tree is constructed by recursively splitting the dataset along the median of a selected coordinate axis until a maximum depth $d=\lfloor\log_{2}(n/s)\rfloor$ is reached.
+Let the input size be $n$, and the time complexity of building the tree be $T(n)$.
+This follows the recurrence: $T(n)=2T(n/2)+\mathcal{O}(n\log_{2}n).$ The recursion terminates after $d$ levels, yielding a total KD-tree construction complexity of $\mathcal{O}(nd\log_{2}(n))$.
+In the second step of geometric compaction, Algorithm 1 is applied independently within each KD-tree block with $s$ Gaussians reduced to $\rho s$ Gaussians.
+Based on our analysis in the previous paragraph, the per-block cost is $O(\rho s^{2})$.
+This along with the fact that there are $2^{d}=n/s$ blocks, the total complexity becomes: $\mathcal{O}\left((n/s)\cdot\rho s^{2}\right)=\mathcal{O}\left(\rho ns\right)={\mathcal{O}}(ms).$ Combining with the KD-tree construction cost, the total time complexity of the geometric compaction step is $\mathcal{O}\left(n(d\log n+mT/{2^{d}})\right).$
+B.3 Detailed Experiments Steps
+To ensure a fair comparison in our experiments, all methods undergo 30,000 total iterations under consistent training conditions.
+MCMC uses 30K iterations of its own update process in [24].
+For other baselines, we summarize the backbone architecture, compaction methods, and fine-tuning steps in the table below.
+Method Backbone Compaction Fine-tune
+(0-15k iterations) (15001th iteration) (150001-30K iterations)
+3DGS a None d
+3DGS+GHAP a Our compaction d
+LightGaussian a LightGaussian compaction d
+PUP-3DGS a PUP-3DGS compaction d
+Trimming the FAT a Trimming the FAT compaction d
+MesonGS a MesonGS compaction d
+MiniSplatting b MiniSplatting compaction d
+MiniSplatting-D+GHAP b b Our compaction d
+LocoGS c LocoGS compaction d
+  • a
+    Vanilla 3DGS densification and pruning in [3].
+  • b
+    Mini-Splatting densification and pruning in [12].
+  • c
+    LocoGS 3DGS update in [33].
+  • d
+    3DGS fine-tuning in [3].
+Key implementation details:
+  • For compaction methods applicable to pre-trained models (e.g., LightGaussian, PUP-3DGS, Trimming the FAT, MesonGS), we initialize from the same backbone model (trained using vanilla 3DGS) and apply their respective compaction directly—excluding any compression modules.
+  • All methods undergo identical fine-tuning (15k iterations) to achieve the target retention ratio.
+This standardized protocol ensures that performance differences stem solely from the methods themselves, eliminating variations due to training procedures.
+All experiments are conducted on a server with 256 GB RAM and a 96-core Intel Xeon Platinum 8255C CPU, and on a workstation equipped with five NVIDIA RTX 3090 GPUs, each with 24 GB of VRAM.
+Appendix C An Additional Comparison Experiments
+As documented in prior work, 3DGS-MCMC [24] reinterpretes the 3D Gaussian Splatting (3DGS) process through the lens of Markov Chain Monte Carlo (MCMC), thereby naturally exploring a broader parameter space.
+In particular, it formulates pruning as a state transition within the MCMC framework and incorporates $L_{1}$ regularization, avoiding abrupt hard-threshold deletions and enabling the natural removal of unnecessary Gaussian elements.
+Given these properties, a further comparative analysis with 3DGS-MCMC is warranted.
+To facilitate a fair comparison, we configured 3DGS-MCMC under two settings: one constrained to 300k Gaussians, and another initialized with 3000k Gaussians followed by compaction using our method to reduce the count to 300k.
+Quantitative results on Tandt & Temples are summarized in the table below.
+[…]
+3DGS-MCMC-300k 0.813 22.925 0.239
+3DGS-MCMC-3000k+GHAP 0.827 22.786 0.209
+As evidenced by the results presented in the table, GHAP combined with 3DGS-MCMC yields superior performance compared to the native 3DGS-MCMC approach constrained to 300k Gaussians.
+This demonstrates the effectiveness of our method in accurately approximating the surface geometry of 3DGS, highlighting the advantage of our compaction strategy.
+Appendix D Ablation on Joint Geometry and Appearance Fine-tuning
+In our configured compaction process, compaction was performed only once at the 15,001st iteration.
+We conducted an ablation study on the timing of compaction to examine the impact of varying compaction frequencies on the final outcome.
+Two alternative compaction strategies were considered: the first involved compacting to 20% at the 15,001st iteration, followed by an additional 50% compaction at the 20,001st iteration; the second strategy applied compaction of 40%, 50%, and 50% at the 15,001st, 20,001st, and 25,001st iterations, respectively.
+All three strategies ultimately resulted in a final retention rate of 10%.
+Experimental results, as presented in the table below, indicate that jointly refining geometry and appearance over multiple stages does not lead to evidently improved performance.
+Therefore, in practical applications, emphasis should be placed on executing appearance optimization for as long as possible, rather than pursuing multi-stage compaction.
+[…]
+15001 0.1 0.817 23.313 0.242
+15001, 20001 0.2,0.5 0.817 23.44 0.247
+15001, 20001, 25001 0.4,0.5,0.5 0.811 23.359 0.255
+Appendix E Additional Numerical Results and Scene Visualizations
+We report a more comprehensive set of results for the comparison experiments in Table LABEL:tab:20, including one different retention ratio: 20%.
+As shown in the updated results, our compaction method significantly outperforms other methods either post-processing compression methods or end-to-end methods.
 In addition to the previously shown Figure 5, we present more detailed qualitative comparisons across multiple scenes in Figure 8.
 Our method consistently preserves the visual quality of the original models.
-Notably, when applied to stronger 3DGS variants with improved densification strategies—such as Mini-Splatting-D—our compaction framework performs even better.
+In particular, when applied to stronger 3DGS variants with improved densification strategies, such as Mini-Splatting-D, our compaction framework performs even better.
 This is reflected in the fact that, after compaction, Mini-Splatting-D often achieves higher rendering quality than the original 3DGS baseline.
-Table 4: Additional information in plug-in experiment.
-Method-Compact Rate $\rho$ Tank&Temples MipNeRF-360 Deep Blendering
-SSIM PSNR LPIPS $k$ Gaussians FPS SSIM PSNR LPIPS $k$ Gaussians FPS SSIM PSNR LPIPS $k$ Gaussians FPS
-3DGS-30k 10% 0.818 23.312 0.242 157 341 0.764 26.404 0.314 272 272 0.905 29.647 0.264 248 290
-20% 0.835 23.615 0.212 313 300 0.788 26.973 0.275 545 230 0.907 29.864 0.252 496 237
-100% 0.853 23.785 0.169 1577 168 0.813 27.554 0.221 2625 126 0.907 29.816 0.238 2476 126
-Mini-Splatting-D 10% 0.835 23.232 0.198 147 564 0.802 27.090 0.25 318 389 0.909 30.042 0.254 156 549
-20% 0.855 23.403 0.171 293 526 0.821 27.310 0.214 636 299 0.912 30.17 0.238 318 489
-100% 0.848 23.338 0.14 1473 209 0.832 27.486 0.176 3177 111 0.907 29.980 0.211 1597 197
-Atom-GS 10% 0.793 22.988 0.274 426 401 0.764 26.535 0.307 468 304 0.899 29.347 0.282 464 419
-20% 0.812 23.282 0.24 852 283 0.788 27.025 0.269 937 236 0.902 29.347 0.268 925 323
-100% 0.814 23.289 0.235 4274 87 0.796 27.135 0.251 4699 86 0.902 29.314 0.267 4625 121
+Table 4: We compared GHAP with four post-processing methods (LightGaussian, PUP-3DGS, Trimming the Fat, MesonGS) at 20% retention, as well as two end-to-end methods (LocoGS, 3DGS-MCMC).
+GHAP also replaces the pruning in Mini-Splatting.
+Results show that our method substantially outperforms both post-processing and end-to-end approaches.
+Method Tanks&Temples MipNeRF-360 Deep Blending
+[…]
+original 3DGS 0.853 23.785 0.169 1577 0.813 27.554 0.221 2627 0.907 29.816 0.238 2475
+LocoGS 0.843 23.655 0.191 571 0.798 27.049 0.257 674 0.903 29.972 0.261 529
+3DGS+GHAP (ours) 0.835 23.615 0.212 314 0.788 26.973 0.275 527 0.907 29.864 0.252 496
+LightGaussian-20% 0.779 22.486 0.271 315 0.765 26.353 0.288 526 0.873 28.011 0.315 495
+PUP-3DGS-20% 0.809 22.603 0.228 315 0.790 26.671 0.257 525 0.905 29.719 0.248 495
+Trimming the Fat-20% 0.819 22.498 0.232 315 0.781 26.494 0.280 524 0.900 29.082 0.272 494
+MesonGS-20% 0.822 20.699 0.207 314 0.776 25.006 0.262 527 0.897 28.696 0.262 496
+3DGS-MCMC 0.779 22.141 0.282 315 0.763 25.957 0.309 263 0.885 28.976 0.298 496
+MiniSplatting-20% 0.824 22.953 0.223 142 0.794 26.728 0.267 215 0.904 29.763 0.265 240
+MiniSplatting+GHAP (ours) 0.855 23.403 0.171 155 0.821 27.310 0.214 219 0.912 30.170 0.238 250
 Figure 8: More scenes visualization. Visual comparisons across additional scenes before and after compaction.